### v1

# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

Abstract
Large language models (LLMs) have been applied in many fields with rapid development in recent years. As a classic machine learning task, time series forecasting has recently received a boost from LLMs. However, there is a research gap in the LLMs’ preferences in this field. In this paper, by comparing LLMs with traditional models, many properties of LLMs in time series prediction are found. For example, our study shows that LLMs excel in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity. We explain our findings through designing prompts to require LLMs to tell the period of the datasets. In addition, the input strategy is investigated, and it is found that incorporating external knowledge and adopting natural language paraphrases positively affects the predictive performance of LLMs for time series. Overall, this study contributes to insight into the advantages and limitations of LLMs in time series forecasting under different conditions.
1 Introduction
Recently, large language models (LLMs) have been widely used, achieving promising performance across various domains, such as health management, customer analysis, and text feature mining Peng et al. (2023); Ledro et al. (2022); Huang et al. (2023). Time series forecasting requires extrapolation from sequential observations. Language models, designed to discern intricate concepts within temporally correlated sequences, intuitively appear well-suited for this task. Hence, LLMs demonstrate proficiency in the domain of time series forecasting Gruver et al. (2023); Rasul et al. (2023); Sun et al. (2023).
Currently, the application of LLMs for time series prediction is at a nascent stage, with the boundaries of this research area remaining ill-defined. The utilization of proprietary large language models in this domain lacks a clear, established methodology. Our objective is to shed light on this emerging field, offering valuable insights and guidance for future research endeavors. However, the LLMs’ preferences for the input time series remain unexplored, such as their proficiency in forecasting both seasonal and trend data. We believe that studying this area clearly will greatly improve the performance of LLM in time series forecasting.
To fill this research gap, in this paper, we focus on the question of what are LLMs’ preferences for the input time series in time series forecasting. To answer this, we conduct experiments on both real and synthesized datasets. Our observations reveal that LLMs perform better on time series characterized by higher trend or seasonal strengths. To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences. We find that LLMs are sensitive to the segment of input sequences proximate to the output.
Naturally, we are interested in the question: Why do LLMs forecast well on datasets with higher trend or seasonal strengths? To solve this, we design prompts that require LLMs to tell the period of the datasets. Through experiments, we let the large language model tell the period of the dataset several times and take the median. We found that the large language model can accurately pinpoint the periodicity of a dataset. This can explain why large language models can predict data sets with high trends or seasonal intensities well since they already learned this kind of knowledge.
In light of these findings, our focus lies on how to leverage these insights to further improve model performance. To address this, we propose two simple techniques to enhance model performance: incorporating external human knowledge and converting numerical sequences into natural language counterparts. Incorporating supplementary information enables large language models to more effectively grasp the periodic nature of time series data, moving beyond a mere emphasis on the tail of the time series. Transforming numerical data into a natural language format enhances the model’s ability to comprehend and reason, also serving as a beneficial approach. Both approaches improve model performance and contribute to our understanding of LLMs in time series forecasting. The workflow is illustrated in Figure 1.
The key contributions are as follows:
We investigate the LLMs’ preferences for the input sequences in time series forecasting. Based on our analysis, we have observed that LLMs are more effective in datasets that have clear periods and trends. A surprising finding is that LLMs may generate output sequences that exhibit higher trend and seasonal strengths. Besides, our results reveal a limitation in LLMs’ performance when confronted with multi-period time series datasets. It indicates that LLMs face challenges in recognizing the distinct periods inherent in such datasets. To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual instances. Our results indicate that large language models are sensitive to the segment of input sequences proximate to the output. All these findings suggest a potential area for improvement in LLMs’ adaptability to complex temporal structures.
Through prompt design and the use of in-context Learning, we have verified that LLMs make better predictions on datasets that exhibit strong trends or seasonality, showing that these models can precisely identify dataset periodicity.
We propose two simple techniques to improve model performance and find that incorporating external knowledge in prompts and paraphrasing natural language positively affects the performance of LLMs in time series forecasting.
2 Preliminaries
2.1 Large Language Model
We use LLMs as a zero-shot learner for time series forecasting by treating numerical values as text sequences. The success of LLMs in time series forecasting can significantly depend on correct pre-processing and handling of the data Gruver et al. (2023). We followed their approach and this process involves a few crucial steps.
In the pre-processing phase for time series forecasting with LLMs, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing. For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space. Tokenization is equally pivotal, shaping the model’s pattern recognition capabilities. Unlike traditional methods like byte-pair encoding (BPE) Hugging Face (2023), which may disrupt numerical coherence, spacing digits ensures individual tokenization, enhancing pattern discernment.
Moreover, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so a specific percentile aligns to 1, facilitating the model’s exposure to varying digit counts and supporting the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
2.2 Time Series Forecasting
In the context of time-series forecasting, the primary goal is to predict the values for the next steps based on observed values from the preceding steps, which is mathematically expressed as:
Here, represents the -step estimation given the previous -step values . denotes the trained parameters from the model , and denotes the prompt or any other information used for the inference. This paper predominantly focuses on single-variate time series forecasting.
Motivated by the interpretability requirements in real-world scenarios, time series can often be decomposed into the trend component, the seasonal component, and the residual component through the addictive model Cleveland et al. (1990). The trend component captures the hidden long-term changes in the data, such as the linear or exponential pattern. The seasonal component captures the repeating variation in the data, and the residual component captures the remaining variation in the data after the removal of trend and seasonal components. This decomposition offers a method to quantify the properties of the time series, which is detailed in section 3.2.
3 What are LLMs’ preferences in time series forecasting?
To explore the preference in the large language model, we first quantify the properties of the input time series to investigate the LLMs’ preferences for time series. Then, to further emphasize our findings, we evaluate the importance of different segments of the input sequence by adding Gaussian noise to the original time series.
3.1 Analyzing Method
To understand the preferences of the LLMs in time series forecasting, we compare LLMTime using various foundational models like GPT-4 and GPT-3.5-turbo traditional methods on various datasets. We also design experiments on synthesized datasets to validate our findings and analyze the impact of the multiple periods. To quantify the LLMs’ preferences towards time series, following Wang et al. (2006), we define the strength of the trend and the seasonality as follows:
where , and denote the trend component, the seasonal component and the residual component respectively. The presented indices serve as indicators of the strength of the trend and seasonality, providing a measure ranging up to 1. It is easy to find that a higher value indicates a stronger trend or seasonality within the time series.
To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual examples. We initiate by defining a sliding window that constitutes 10% of the total length of the time series. Given the variability in the length of time sequences, we first scale the sequence before incorporating the noise into the original sequence. This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLMs predominantly focus on.
3.2 Preferences for the time series Experiment
In this subsection, we delve into the input sequence preferences for time series forecasting with LLMs. Our experiments are conducted on both real datasets and synthetic datasets. Throughout this subsection, we use GPT-3.5-turbo-instruct and GPT-4 for time series forecasting and measure model performance through and Mean Absolute Percentage Error (MAPE).
3.2.1 Implementation Details
Real Datasets: We conduct experiments on ten real datasets, as enumerated in Appendix A.2. we apply the Seasonal-Trend decomposition using the LOESS (STL) techniqueCleveland et al. (1990), to decompose the original time series into trend, seasonal, and residual components. Subsequently, we computed the strengths of the trend strength and seasonal strength . To further discern the LLMs’ preferences for the specific segments of the input data, we conduct the counterfactual analysis with a systematic permutation to the input time series. We initiate by defining a sliding window that constitutes 10% of the total length of the time series and add Gaussian noise into the data within this window data. Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence. Refer to A.2.1 for detailed information.
Synthesized Datasets: To further validate our findings, we conduct experiments on synthesized datasets. We formulate the dataset through the expression . To investigate the influence of the number of periods on model performance, we generated a dataset using the function . ranges from 0 to 20 and follows the normal distribution . Refer to A.2.1 for detailed information.
3.2.2 Key Findings
After computing the Pearson correlation coefficients (PCC), we observed a nearly strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series has a higher trend and seasonal strength. Notably, GPT-4 achieved a higher PCC compared to GPT-3.5-turbo-instruct. It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data. Interestingly, there is an increase in the of the output generated by GPT-4, compared to the original test sequences. It indicates that GPT-4 tends to forecast time series with high seasonal strength, which may provide insights for further research. In the context of multi-period time series, the model performance decreases as the number of periods increases. It indicates that LLMs may have trouble recognizing the multiple periods inherent in such datasets, which is common in reality. For the counterfactual analysis, as shown in Figure 6, there is a noticeable decrease in values when Gaussian noise is added to the latter segments. Our findings reveal that LLMs are more sensitive to the end of input time series data when it acts as a time series predictor.
4 Why do LLMs forecast well on datasets with higher trend or seasonal strengths?
Our findings indicate that LLMs exhibit enhanced performance on time series data characterized by pronounced trends or significant seasonal variations. This raises the question: Why do LLMs excel in forecasting datasets with marked trend or seasonal patterns? To explore this phenomenon, we crafted prompts that necessitate LLMs to acknowledge the temporal span of the dataset.
This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct trend or seasonal attributes in-context learningvanced pattern recognition capabilities. By explicitly prompting LLMs to consider the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, thereby shedding light on the mechanisms that underpin their superior performance in such contexts.
4.1 Implementation Details
In order to explore the phenomenon that LLMs forecast well on datasets with higher trend or seasonal strengths, we design experiments to verify this phenomenon. We design prompts to let the LLMs output the predicted value after each sequence prediction. The target model of the experiment is GPT-3.5-Turbo (gpt-3.5-turbo-1106), and the main role of our prompt is to perform in-context learning and require output cycles, as detailed in the Appendix A.3. We select eight time series data sets such as AirPassengersDataset, count the period value after ten times of prediction, and compare the median of these ten results with the real period. The results are shown in Table 3.
| Dataset Name | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| IstanbulTrafficGPT | ||||||
| GasRateCO2Dataset | ||||||
| HeartRateDataset | ||||||
| TurkeyPower |
4.2 Key Findings
According to the results, we find that large language models can determine the periodicity of a dataset to some extent. Although the fluctuation of each time series prediction is large, the prediction on AirPassenger, Sunspots and Wooly datasets is relatively accurate, and the predicted values on other datasets are also close to our real cycle value. We speculate that LLMs can predict datasets with high trend or seasonal intensity well because they already possess some knowledge about the scene and content of the dataset.
5 How to leverage these insights to improve model’s performance?
Based on these findings, our focus is on how to leverage these findings to further improve model performance. Because our ultimate goal is to improve the time series prediction performance of the model. Therefore, we retain the original structure of the time series data but enhance the model’s input with prompts infused with additional knowledge. These prompts are crafted with seasonal trends and cyclical behaviours in the dataset to endow the model with a richer contextual understanding. In another way, we transform time series data into formats that resemble natural language sequences, enabling large language models to leverage their superior inference skills more effectively in the realm of time series prediction.
5.1 External Knowledge Enhancing Time Series Forecasting
We introduce a novel method to improve the performance and stability of large language models for time series forecasting. The core idea of this part is to use the knowledge obtained from pre-training of large language models to help predict. We will give the large language model some basic information about the current dataset such as the name of the dataset in the prompt, and this process does not involve data leakage (it does not tell the large language model about the period or predicted value, etc.) Assuming we define initial prompt as the original time sequence, the extra information defined as , the new prompt can be written as:
5.1.1 Implementation Details
We input the dataset’s external knowledge through prompts before the sequence’s input. The external knowledge of each dataset is presented in Appendix A.1. The external knowledge includes the description of dataset features and scenarios and the dataset’s data properties. The results are shown in Table 5.
5.1.2 Key Findings
In External Knowledge Enhancement, GPT-4 generally performs better than GPT-3.5 on MSE, MAE and MAPE, especially on AirPassengers, AusBeer and other datasets. Llama-2 significantly outperforms GPT-3.5 and GPT-4 in terms of MSE and MAE metrics on some datasets (e.g., Wooly, ETTh1, ETTm2), indicating that it can capture data features more accurately. R-squared values are shown on some datasets such as ETTh1, ETTm2, and Turkey Power, all models can provide relatively accurate predictions with R-squared values close to 1.
In the LLMTime Prediction, GPT-4 and Llama-2 perform relatively well on AirPassengers and AusBeer datasets with R values close to or above 0.5. GPT-3.5 achieves very high MSE on Sunspots and Wine datasets, which may be due to the complexity of the dataset or the poor adaptation of the model for these specific tasks. The Turkey Power dataset has a high R value on the time prediction task for all models, indicating a good model fit for this task.
5.2 Natural Language Paraphrasing of Sequences
We will be creating experiments related to natural language paraphrasing of sequences. This will be based on the fact that LLMs are not affected by the order of magnitude and size of digits Shah et al. (2023).
We have created a method to convert sequences into natural language paraphrases, the main process is using traditional natural language processing techniques to preprocess the series. For instance, if we have a discrete time series that represents temperature changes, we can denote it as , where . To do this, we extract a natural language description of the sequence based on two data points and a trend from to , "This is a discrete temperature time series. The temperature rises from to , and falls from to …". The string we get is our natural language paraphrasing sequence.
The sequence generated by the natural language paraphrasing of the text in is represented by . Finally, we paraphrase the sequence from to through the large language model; we let the LLM prediction of get , after Extraction and Processing, get , then the above completed our prediction task.
We load in the sequence and determine the movement of the sequence at each step. Then add "from xx increasing to xx," or "from xx decreasing to xx." Complete the process of modifying from time series to natural language. To convert a text sequence into a discrete sequence, we use the Reverse function, which converts the natural language series into a regular expression format through hard code.
5.2.1 Implementation Details
We use the LLMTime model with GPT-3.5-Turbo, GPT-4-turbo and Llama-2 as the core to forecast the built-in time series data set, the training set and test set ratio is 4:1. We will get the time series prediction results and the original series MSE, MAE, MAPE and R. The result is present in Table 4 compared with LLMTime forecasting results.
5.2.2 Key Findings
According to the experimental data obtained in the above table, we find that enhancing LLM through natural language paraphrasing has a certain improvement effect on time series analysis and prediction on some datasets. From the perspective of model performance comparison, the R value of GPT-3.5-Turbo in Natural Language Paraphrasing and LLMTime Prediction method is negative in some cases, indicating that the model performs poorly. GPT-4-Turbo performs better on most datasets compared to other magic methods, especially on Natural Language Paraphrasing methods, MSE, MAE, and MAPE usually decrease, while R improves. Llama-2 shows better performance on some datasets, such as lower MSE and MAE and better MAPE and R in the Natural Language Paraphrasing method of the AusBeer dataset.
6 Related Works
6.1 Time Series Prediction
Time series prediction is an important field in the field of artificial intelligence. The unified idea of time series prediction is to analyze past and future series trends. It is used in finance and business to predict the trend of indexes and stocks, in the field of intelligent manufacturing to predict anomaly detection and power load, and in the field of meteorology, agriculture, and navigation to predict temperature, humidity, and climate Singh and Malhotra (2023); Munir et al. (2018); Shen et al. (2020); Braei and Wagner (2020). We categorize related research within this domain into three categories: traditional time series prediction, time series prediction based on machine learning, and deep learning.
| Metrics | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
| GPT4-MAPE | ||||||
| GPT4-R2 | ||||||
| GPT3.5-MAPE | ||||||
| GPT3.5-R2 | ||||||
| Trend Strength | ||||||
| Seasonal Strength |
Two commonly used methods for traditional time series analysis are the ARIMA method Box and Pierce (1970) and the exponential smoothing method Gardner Jr (2006). The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions. On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations. The ARIMA model requires testing the stationarity of data and selecting the right order. On the other hand, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower compared to the ARIMA model.
| Dataset | Period | Real | Median | |||||||||
| AirPassengersDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
| AusBeerDataset | 24 | 10 | 6 | 6 | 1 | 3 | 6 | 7 | 3 | 21 | 4 | 6 |
| GasRateCO2Dataset | 7 | 36 | 24 | 72 | 28 | 34 | 20 | 7 | 16 | 14 | 39 | 20 |
| MonthlyMilkDataset | 6 | 6 | 25 | 11 | 6 | 2 | 7 | 34 | 8 | 7 | 12 | 7 |
| SunspotsDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
| WineDataset | 1 | 11 | 12 | 11 | 12 | 7 | 6 | 8 | 10 | 12 | 12 | 10 |
| WoolyDataset | 10 | 5 | 6 | 7 | 16 | 11 | 5 | 11 | 3 | 4 | 4 | 5 |
| HeartRateDataset | 13 | 16 | 6 | 12 | 13 | 11 | 12 | 32 | 32 | 64 | 144 | 12 |
Time series analysis based on linear regression, decision tree random forest, and other methods can make predictions for future series through the analysis and learning of existing series, as more complex time series prediction methods based on deep learning, RNN, CNN, and LSTM have played a huge role in many applications. Especially in the field of sequence prediction, many variants of LSTM have emerged, such as bidirectional LSTM Graves et al. (2005), gated recurrent unit Chung et al. (2014), long short-term memory network Malhotra et al. (2015), and deep recurrent neural network Pascanu et al. (2013). This method improves prediction accuracy but takes longer to train and is less prone to overfitting than the original LSTM.
6.2 LLMs for Time Series Prediction
Given the notable achievements of LLMs across various domains, it is prudent to incorporate time series prediction tasks with LLMs. Similar to (Sun et al., 2023), we categorize contemporary research within this domain into two categories: LLM-for-Time Series and Time Series-for-LLM.
LLM-for-Time Series involves either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data Rasul et al. (2023); Garza and Mergenthaler-Canseco (2023); Das et al. (2023); Cao et al. (2023). For instance, (Rasul et al., 2023) aimed to build the foundational models for time series and investigate its scaling behavior. Chang et al. (2024) proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting. Although these studies contribute significantly to understanding foundational models, they demand considerable computing resources and expertise in fine-tuning procedures. Moreover, the details of the model may not be disclosed for commercial purposes Garza and Mergenthaler-Canseco (2023), which impedes future research. Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning. Therefore, in this paper, we concentrate on the Time Series-foR-LLM, harnessing fixed LLMs for the time-series forecasting task.
| Models | Datasets | Natural Language Paraphrasing | LLMTime Prediction | ||||
| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
| GPT-3.5-Turbo | AirPassengers | 267.66 | 3.66 | 0.99 | 6244.07 | 61.39 | 14.43 |
| AusBeer | 598.45 | 5.81 | 1.36 | 841.68 | 23.59 | 5.62 | |
| GasRateCO2 | 3.16 | 0.46 | 0.85 | 10.88 | 2.66 | 4.73 | |
| MonthlyMilk | 968.69 | 8.61 | 1.02 | 7507.13 | 66.28 | 112.77 | |
| Sunspots | 251.61 | 4.27 | 20.42 | 6556.55 | 58.95 | 217.94 | |
| Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 | |
| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
| GPT-4-Turbo | AirPassengers | 133.10 | 2.87 | 0.80 | 1286.25 | 28.04 | 6.07 |
| AusBeer | 661.80 | 7.24 | 1.63 | 513.49 | 18.57 | 4.28 | |
| GasRateCO2 | 2.28 | 0.41 | 0.75 | 7.27 | 2.32 | 4.18 | |
| MonthlyMilk | 413.63 | 4.94 | 0.57 | 4442.18 | 50.75 | 172.82 | |
| Sunspots | 194.52 | 5.30 | 16.10 | 3374.70 | 41.87 | 321.11 | |
| Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 | |
| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1286.25 | 28.04 | 6.07 |
| AusBeer | 591.75 | 23.25 | 5.41 | 644.82 | 17.88 | 4.08 | |
| GasRateCO2 | 10.16 | 2.89 | 5.16 | 12.78 | 2.97 | 5.47 | |
| MonthlyMilk | 851.17 | 84.83 | 9.46 | 3410.20 | 41.40 | 240.25 | |
| Sunspots | 1483.29 | 33.27 | 17.79 | 4467.67 | 48.95 | 91.79 | |
| Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 | |
| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
Time Series-for-LLM centres around using existing LLMs and designing mechanisms, such as crafting appropriate prompts or reprogramming inputs, to handle time series data effectively (Gruver et al., 2023; Sun et al., 2023; Jin et al., 2023; Xue and Salim, 2023). Sun et al. (2023) tokenized the time series and managed to embed those tokens, and Jin et al. (2023) reprogrammed the time series data with text prototypes before feeding them to the LLMs. These studies illuminated the characteristics of time series data and devised methods to align them with LLMs. However, they lack an analysis of the ability and bias in forecasting time series. The most related work to us is Gruver et al. (2023), though, it lacks a quantitative analysis of the preference for the time series in LLMs, and it failed to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt. Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
7 Conclusions
Our research has unearthed key preferences of large language models (LLMs) in the domain of time series forecasting, revealing a proclivity for data with distinct trends and seasonal patterns. Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity. Our proposed strategies incorporating external knowledge and transforming numbers into natural language have both enhanced LLMs’ predictive accuracy. Despite encountering challenges with multi-period time series, our work provides a pivotal foundation for future advancements, offering actionable insights and methods to refine the application of LLMs in time series analysis.
Limitation
This study may be limited in the following ways. First, limitations in the scope of the dataset and large language models may not capture the full variability of the results with a wider array. In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods. Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the performance of the model in different data domains. These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, which underscores the need to expand future research.
Appendix A Appendix
A.1 The External Knowledge incorporated in the Prompt
(1) AirPassengersDataset: This is a series of monthly passenger numbers for international flights, where each value is in thousands of passengers for that month.
(2) AusBeerDataset: This is a quarterly series of beer production, with each value representing the kilolitres of beer produced in that quarter.
(3) GasRateCO2Dataset: This time series dataset describes monthly carbon dioxide emissions.
(4) MonthlyMilkDataset: This is a time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
(5) SunspotsDataset: This is a dataset that records the number of sunspots in each month, where each data is the number of sunspots in that month.
(6) WineDataset: This is a dataset of monthly wine production in Australia, where each figure is the number of wine bottles produced in that month.
(7) WoolyDataset: This is an Australian yarn production for each quarter, where each value is how many tons of yarn were produced in that quarter.
(8) HeartRateDataset: The series contains 1800 uniformly spaced instantaneous heart rate measurements from a single subject.
(9) Weather: This dataset contains weather-related data such as temperature, precipitation, humidity, etc. It is commonly used in weather forecasting and climate research.
(10) COVID Deaths: This is a collection of data documenting related deaths during the COVID-19 pandemic, including information on the number of deaths, infection rates, etc., for use in outbreak analysis and health policy development.
(11) Solar Weekly: This dataset contains weekly data on solar power generation, which can be used to analyze and forecast renewable energy production.
(12) Tourism Monthly: This dataset records monthly tourism-related data such as the number of tourists, tourism revenue, etc., for tourism trend analysis and prediction.
(13) Australian Electricity Demand: This deals with electricity demand data in Australia, which can be used for electricity market analysis and grid management.
(14) Pedestrian Counts: Pedestrian count data are recorded for specific areas, such as city streets, for urban planning and traffic management.
(15) Traffic Hourly: Provides hourly traffic flow data for traffic forecasting and urban planning.
(16) Hospital: This dataset may contain hospital-related data such as admission rates, bed occupancy rates, etc., for medical resource planning and public health research.
(17) Fred MD: It is probably a collection of economic data, such as gross domestic product (GDP), unemployment rate, etc., used for macroeconomic analysis and forecasting.
(18) Tourism Yearly: Similar to "Tourism Monthly", but provides annual tourism data for longer-term analysis of tourism trends.
(19) Tourism Quarterly: Provides quarterly tourism data between monthly and annual data for medium-term tourism trend analysis.
(20) US Births: A dataset that records birth rates in the United States, including information such as the number of births, for use in population research and social policy making.
(21) NN5 Weekly: This may be a weekly financial or economic dataset, including stock market data, economic indicators, and so on.
(22) Traffic Weekly: Provides weekly traffic flow data for long-term traffic pattern analysis and planning.
(23) Saugeenday: This may be a daily degree dataset for a specific region, such as the Saugeen region, and the data involved may include environmental, meteorological, or socioeconomic data.
(24) CIF 2016: This may be a specific economic or financial time series dataset collected in 2016 for economic forecasting and analysis.
(25) Bitcoin: This dataset deals with the price and transaction data of Bitcoin and is used for cryptocurrency market analysis and forecasting.
(26) Sunspot: Data recording sunspot activity for astronomical research and prediction of the solar cycle.
(27) NN5 Daily: This is similar to NN5 Weekly, but provides daily data that may refer to financial markets or economic indicators.
A.2 Details of the experiments on the LLMs’ Preferences
In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data. We first describe both the real and synthesized datasets we use, and then detail the methods we use to investigate the preferences of LLMs.
A.2.1 Real Datasets
We conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTraffic and TurkeyPower, as detailed in A.1. We apply the Seasonal-Trend decomposition using the LOESS (STL) technique, to decompose the original time series into trend, seasonal, and residual components. In those datasets, we obtain the periods through the nature of the data. For instance, the number of passengers is collected every month in AirPassengersDataset, and it’s natural to obtain that the period is 12. For the datasets without explicit periods, such as the IstanbulTraffic, the period was determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period. The strengths and the model performance can be seen in Table 1.
Subsequently, we compute trend strength and seasonal strength to measure all of those components. We use and MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 2). Notably, GPT-4 achieved a higher absolute PCC compared to GPT-3.5-turbo-instruct. It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data. Interestingly, upon calculating the strengths of the output sequences, there is an increase in the seasonal component of the output generated by GPT-4, in contrast to the original test sequences. This may provide some insights for further research into the characteristics of the LLMs.
Following this, we conduct the counterfactual analysis with a systematic permutation to the input time series. We initiate by defining a sliding window that constitutes 10% of the total length of the time seriesm and add Gaussian noise into the data within this window data. Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence. The size of the sliding window, , is defined as: , where is the total length of the time series. We denote the start point of the window as and the Gaussian noise as , to the data within the sliding window, the modified data point, , is given by:
where , and is the data point after normalization. The change in Mean Squared Error () is defined as the metric to measure the sensitivity of the LLMs:
where as , reaches its maximum.
This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLM predominantly focuses on. Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
A.2.2 Synthesized Datasets
To further validate our findings, we conducted experiments on synthesized datasets. We formulate the dataset through the expression , where ranges from 0 to 20, and the noise term follows a normal distribution . The coefficients and were uniformly sampled for 10 samples respectively, with and . Hence, both trend and seasonal strengths were carefully confined to ,, resulting in more discernible conclusions in the output. The chosen frequency, , was consistent across the experiments. We uniformly sampled 200 data points. Our observations validate that LLMs prefer input sequences with higher strengths and generate time series with higher seasonal strength. Furthermore, we observed that the model performance cannot be entirely explained by the linear combination of the strengths, implying that the influence of these strengths is not independent.
To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components. We set , , and , uniformly sampled for 10 instances each, and and to represent the chosen frequencies. Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution . Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure6(a). This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
A.3 Detail Prompts
Request output periodic prompts
Please output the period size of the predicted sequence as an integer. Output an integer directly.
In-context learning examples
For example, the sequence: 0.387952, 8.975192, 5.398713, -6.011139, -9.807413, -0.261663, 9.245404, 5.901009, -6.038525, -9.966693, -0.022652, 9.352823,5.317264, -5.809847, -9.565386, -0.325205, 8.711379, 6.176601, -5.911722, -9.078278.We can get the period number is 5. so, try to find out this period:
A.4 Figures and Tables
The predicted results of AirPassengers, AusBeerDataset, GasRateCO2, HeartRate, Istanbul-Traffic datasets.
The predicted results of MonthlyMilk, Sunspots, TSMCStock, TurkeyPower, WineDataset, Wooly datasets.
| Models | Dataset | External Knowledge Enhancing | LLMTime Prediction | ||||
| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
| AirPassengers | 3713.99 | 50.37 | 10.88 | 6244.07 | 61.39 | 14.43 | |
| AusBeer | 669.01 | 21.82 | 5.12 | 707.43 | 20.44 | 4.80 | |
| GasRateCO2 | 16.47 | 3.36 | 5.97 | 10.88 | 2.66 | 4.73 | |
| HeartRate | 59.83 | 6.44 | 6.75 | 3863.65 | 48.90 | 5.49 | |
| MonthlyMilk | 4781.26 | 55.45 | 6.25 | 7507.13 | 66.28 | 112.77 | |
| Sunspots | 7072.42 | 62.61 | 194.29 | 33094824.12 | 4209.81 | 17.69 | |
| GPT-3.5 | Wine | 24925885.81 | 3548.19 | 14.98 | 820422.05 | 760.03 | 16.51 |
| (gpt-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 65.95 | 7.05 | 7.47 |
| Istanbul-Traffic | 888.31 | 28.16 | 60.11 | 1321.44 | 48.7 | 7.47 | |
| TSMC-Stock | 73.83 | 7.31 | 1.54 | 298.58 | 15.44 | 3.23 | |
| Turkey Power | 2613198.17 | 1301.83 | 3.52 | 3882704.14 | 1315.6 | 3.58 | |
| ETTh1 | 2.65 | 1.01 | 132.13 | 5.64 | 2.71 | 1.625 | |
| ETTm2 | 2.00 | 0.89 | 201.84 | 3.46 | 2.17 | 1.178 | |
| Wind | 36.05 | 4.99 | inf | 32.07 | 3.76 | 10.41 | |
| AirPassengers | 1262.24 | 30.54 | 6.80 | 1286.25 | 28.04 | 6.07 | |
| AusBeer | 345.59 | 15.70 | 3.69 | 513.49 | 18.57 | 4.28 | |
| GasRateCO2 | 6.99 | 2.29 | 4.21 | 7.27 | 2.32 | 4.18 | |
| HeartRate | 78.99 | 6.96 | 7.90 | 1878.24 | 40.79 | 4.74 | |
| MonthlyMilk | 2209.33 | 44.02 | 5.12 | 4442.18 | 50.75 | 172.82 | |
| Sunspots | 4571.92 | 50.24 | 334.30 | 10105667.50 | 2438.74 | 9.85 | |
| GPT-4 | Wine | 14426570.88 | 2734.41 | 10.90 | 1211754.67 | 1017.44 | 21.85 |
| (gpt-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 85.22 | 6.89 | 7.77 |
| Istanbul-Traffic | 954.88 | 26.92 | 47.29 | 1291.17 | 32.16 | 6.46 | |
| TSMC-Stock | 104.53 | 8.46 | 1.79 | 74.71 | 6.60 | 1.39 | |
| Turkey Power | 3090055.89 | 1223.78 | 3.36 | 113873.28 | 814.46 | 2.17 | |
| ETTh1 | 2.70 | 1.06 | 129.99 | 4.73 | 1.53 | 3.282 | |
| ETTm2 | 1.18 | 0.79 | 291.67 | 2.30 | 1.034 | 1.607 | |
| Wind | 1.82 | 0.81 | 82.50 | 28.83 | 6.32 | 2.14 | |
| AirPassengers | 3713.99 | 50.37 | 10.88 | 1286.25 | 28.04 | 6.07 | |
| AusBeer | 893.56 | 21.49 | 4.87 | 644.82 | 17.88 | 4.08 | |
| GasRateCO2 | 11.38 | 3.04 | 5.49 | 12.78 | 2.97 | 5.47 | |
| HeartRate | 112.17 | 7.86 | 8.93 | 3317.83 | 52.71 | 6.15 | |
| MonthlyMilk | 4722.32 | 60.36 | 7.05 | 3410.20 | 41.40 | 240.25 | |
| Sunspots | 4000.19 | 46.45 | 138.69 | 11788774.81 | 2770.67 | 10.90 | |
| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 636637.46 | 742.67 | 15.54 |
| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 64.92 | 6.39 | 7.04 |
| Istanbul-Traffic | 979.15 | 26.70 | 45.57 | 1531.37 | 34.74 | 7.42 | |
| TSMC-Stock | 52105.36 | 196.02 | 42.07 | 2203.97 | 27.64 | 27.39 | |
| Turkey Power | 3416162.71 | 1547.49 | 4.09 | 2919773.15 | 1388.10 | 3.70 | |
| ETTh1 | 4.15 | 1.65 | 408.11 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 3.08 | 1.47 | 810.56 | 3.31 | 2.07 | 2.153 | |
| Wind | 1.78 | 0.79 | 116.31 | 36.78 | 4.02 | 12.38 | |
| Dataset | Method | MSE | MAE | MAPE |
| AirPassengers | Exponential Smoothing | 2007.67 | 37.91 | 8.10 |
| SARIMA | 2320.47 | 39.80 | 8.46 | |
| Cyclical Regression | 2028.37 | 36.70 | 8.52 | |
| AutoARIMA | 8702.09 | 68.52 | 13.98 | |
| FFT | 3274.46 | 46.38 | 10.59 | |
| StatsForecastAutoARIMA | 2952.52 | 45.41 | 9.71 | |
| Naive Mean | 47703.65 | 204.25 | 44.61 | |
| Naive Seasonal | 6032.80 | 62.87 | 14.18 | |
| Naive Drift | 6505.79 | 72.21 | 17.50 | |
| Naive Moving Average | 6032.80 | 62.87 | 14.18 | |
| LLMTime with GPT-3.5-Turbo | 267.66 | 3.66 | 0.99 | |
| LLMTime with GPT-4-Turbo | 133.10 | 2.87 | 0.80 | |
| AusBeer | Exponential Smoothing | 703.26 | 22.80 | 5.44 |
| SARIMA | 475.53 | 19.07 | 4.49 | |
| Cyclical Regression | 989.31 | 26.29 | 6.13 | |
| AutoARIMA | 550.05 | 18.84 | 4.41 | |
| FFT | 7682.56 | 73.74 | 17.44 | |
| StatsForecastAutoARIMA | 559.46 | 20.56 | 4.86 | |
| Naive Mean | 1885.72 | 30.66 | 6.68 | |
| Naive Seasonal | 10828.02 | 96.35 | 23.39 | |
| Naive Drift | 18507.61 | 128.23 | 30.91 | |
| Naive Moving Average | 10828.02 | 96.35 | 23.39 | |
| LLMTime with GPT-3.5-Turbo | 598.45 | 5.81 | 1.36 | |
| LLMTime with GPT-4-Turbo | 661.80 | 7.24 | 1.63 | |
| MonthlyMilk | Exponential Smoothing | 564.94 | 20.23 | 2.41 |
| SARIMA | 1289.76 | 32.78 | 3.87 | |
| Cyclical Regression | 3631.53 | 56.15 | 6.60 | |
| AutoARIMA | 2682.67 | 42.82 | 5.20 | |
| FFT | 3453.96 | 45.62 | 5.48 | |
| StatsForecastAutoARIMA | 186.14 | 10.64 | 1.28 | |
| Naive Mean | 19893.07 | 127.33 | 14.46 | |
| Naive Seasonal | 4870.40 | 56.00 | 6.31 | |
| Naive Drift | 3998.11 | 56.06 | 6.52 | |
| Naive Moving Average | 4870.40 | 56.00 | 6.31 | |
| LLMTime with GPT-3.5-Turbo | 968.69 | 8.61 | 1.02 | |
| LLMTime with GPT-4-Turbo | 413.63 | 4.94 | 0.57 | |
| Sunspots | Moving Average | 326750.49 | 499.78 | 3129.63 |
| Exponential Smoothing | 326750.49 | 499.78 | 3129.63 | |
| SARIMA | 2902.72 | 45.75 | 466.99 | |
| Cyclical Regression | 3917.76 | 47.84 | 274.31 | |
| AutoARIMA | 4695.67 | 58.47 | 709.23 | |
| FFT | 3784.56 | 49.81 | 150.32 | |
| StatsForecastAutoARIMA | 8406.55 | 72.99 | 95.18 | |
| Naive Mean | 4120.40 | 49.84 | 267.22 | |
| Naive Seasonal | 4440.63 | 56.78 | 688.58 | |
| Naive Drift | 5032.77 | 60.40 | 724.88 | |
| Naive Moving Average | 4440.63 | 56.78 | 688.58 | |
| LLMTime with GPT-3.5-Turbo | 251.61 | 4.27 | 20.42 | |
| LLMTime with GPT-4-Turbo | 194.52 | 5.30 | 16.10 |
| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| GasRateCO2Dataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| HeartRateDataset | ||||||
| IstanbulTraffic | ||||||
| TurkeyPower |
| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| GasRateCO2Dataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| HeartRateDataset | ||||||
| IstanbulTraffic | ||||||
| TurkeyPower |

### v3

@@
 Abstract
-Large language models (LLMs) have been applied in many fields with rapid development in recent years.
-As a classic machine learning task, time series forecasting has recently received a boost from LLMs.
-However, there is a research gap in the LLMs’ preferences in this field.
-In this paper, by comparing LLMs with traditional models, many properties of LLMs in time series prediction are found.
-For example, our study shows that LLMs excel in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity.
-We explain our findings through designing prompts to require LLMs to tell the period of the datasets.
-In addition, the input strategy is investigated, and it is found that incorporating external knowledge and adopting natural language paraphrases positively affects the predictive performance of LLMs for time series.
-Overall, this study contributes to insight into the advantages and limitations of LLMs in time series forecasting under different conditions.
+Large language models (LLMs) have been applied in many fields and have developed rapidly in recent years.
+As a classic machine learning task, time series forecasting has recently been boosted by LLMs.
+Recent works treat large language models as zero-shot time series reasoners without further fine-tuning, which achieves remarkable performance.
+However, there are some unexplored research problems when applying LLMs for time series forecasting under the zero-shot setting.
+For instance, the LLMs’ preferences for the input time series are less understood.
+In this paper, by comparing LLMs with traditional time series forecasting models, we observe many interesting properties of LLMs in the context of time series forecasting.
+First, our study shows that LLMs perform well in predicting time series with clear patterns and trends but face challenges with datasets lacking periodicity.
+This observation can be explained by the ability of LLMs to recognize the underlying period within datasets, which is supported by our experiments.
+In addition, the input strategy is investigated and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series.
+Overall, our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
 1 Introduction
-Recently, large language models (LLMs) have been widely used, achieving promising performance across various domains, such as health management, customer analysis, and text feature mining Peng et al.
-(2023); Ledro et al.
-(2022); Huang et al.
-(2023).
+Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [16, 15, 12].
 Time series forecasting requires extrapolation from sequential observations.
-Language models, designed to discern intricate concepts within temporally correlated sequences, intuitively appear well-suited for this task.
-Hence, LLMs demonstrate proficiency in the domain of time series forecasting Gruver et al.
-(2023); Rasul et al.
-(2023); Sun et al.
-(2023).
-Currently, the application of LLMs for time series prediction is at a nascent stage, with the boundaries of this research area remaining ill-defined.
-The utilization of proprietary large language models in this domain lacks a clear, established methodology.
-Our objective is to shed light on this emerging field, offering valuable insights and guidance for future research endeavors.
-However, the LLMs’ preferences for the input time series remain unexplored, such as their proficiency in forecasting both seasonal and trend data.
-We believe that studying this area clearly will greatly improve the performance of LLM in time series forecasting.
-To fill this research gap, in this paper, we focus on the question of what are LLMs’ preferences for the input time series in time series forecasting.
-To answer this, we conduct experiments on both real and synthesized datasets.
-Our observations reveal that LLMs perform better on time series characterized by higher trend or seasonal strengths.
+Language models are designed to discern intricate concepts within temporally correlated sequences, and intuitively appear well-suited for this task.
+Hence, there exist some preliminary studies that apply LLMs to time series forecasting tasks [10, 17, 19].
+However, currently the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined.
+There are many unexplored problems in this field.
+For example, existing research lacks exploration into how the performance of LLMs varies when faced with different types of time series inputs.
+This includes the effectiveness gap for LLMs in predicting data with seasonal and trending patterns versus data without such patterns.
+To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero shot prompting setting.
+Through experiments on both real and synthesized datasets, we find that LLMs perform better in time series with higher trend or seasonal strengths.
+Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture the distinct periods within those datasets.
 To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences.
-We find that LLMs are sensitive to the segment of input sequences proximate to the output.
-Naturally, we are interested in the question: Why do LLMs forecast well on datasets with higher trend or seasonal strengths?
-To solve this, we design prompts that require LLMs to tell the period of the datasets.
-Through experiments, we let the large language model tell the period of the dataset several times and take the median.
-We found that the large language model can accurately pinpoint the periodicity of a dataset.
-This can explain why large language models can predict data sets with high trends or seasonal intensities well since they already learned this kind of knowledge.
-In light of these findings, our focus lies on how to leverage these insights to further improve model performance.
+The findings suggest that LLMs are particularly sensitive to the segment of input sequences closest to the target output.
+Based on the above findings, we want to further explore why LLMs forecast well on datasets with higher seasonal strengths.
+To this end, we require LLMs to tell the period of the datasets through multiple runs.
+We find that LLMs can mostly recognize the underlying period of a dataset.
+This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well, since they can obtain the seasonal pattern inside the datasets.
+In light of the above-mentioned findings, we are interested in how to leverage these insights to further improve model performance.
 To address this, we propose two simple techniques to enhance model performance: incorporating external human knowledge and converting numerical sequences into natural language counterparts.
@@
 The key contributions are as follows:
-We investigate the LLMs’ preferences for the input sequences in time series forecasting.
-Based on our analysis, we have observed that LLMs are more effective in datasets that have clear periods and trends.
-A surprising finding is that LLMs may generate output sequences that exhibit higher trend and seasonal strengths.
-Besides, our results reveal a limitation in LLMs’ performance when confronted with multi-period time series datasets.
-It indicates that LLMs face challenges in recognizing the distinct periods inherent in such datasets.
-To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual instances.
-Our results indicate that large language models are sensitive to the segment of input sequences proximate to the output.
-All these findings suggest a potential area for improvement in LLMs’ adaptability to complex temporal structures.
-Through prompt design and the use of in-context Learning, we have verified that LLMs make better predictions on datasets that exhibit strong trends or seasonality, showing that these models can precisely identify dataset periodicity.
-We propose two simple techniques to improve model performance and find that incorporating external knowledge in prompts and paraphrasing natural language positively affects the performance of LLMs in time series forecasting.
+We investigate the preferences for the input sequences in LLMs in time series forecasting tasks.
+Our analysis has revealed that LLMs significantly outperform traditional time series forecasting methods without the need for additional fine-tuning.
+Interestingly, LLMs display superior predictive capabilities when dealing with datasets that have higher trends and seasonal strengths.
+We require LLMs to identify the periodicity of datasets across multiple iterations.
+Our observations indicate that LLMs can effectively recognize the inherent periodic patterns within datasets.
+This observation answers the question of why LLMs perform well in forecasting time series with higher seasonal strengths, as they can capture the seasonal patterns inherent in the data.
+We propose two simple techniques to improve model performance and find that both incorporating external human knowledge into input prompts and paraphrasing input sequences to natural language substantially improve the performance of LLMs in time series forecasting.
 2 Preliminaries
@@
 We use LLMs as a zero-shot learner for time series forecasting by treating numerical values as text sequences.
-The success of LLMs in time series forecasting can significantly depend on correct pre-processing and handling of the data Gruver et al.
-(2023).
-We followed their approach and this process involves a few crucial steps.
-In the pre-processing phase for time series forecasting with LLMs, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing.
+In this paper, we investigate three close source LLMs, including GPT-3.5-turbo, GPT-4-turbo, and Gemini-1.0-Pro, and one open-source LLMs, i.e., llama-2-13B.
+The success of LLMs in time series forecasting can significantly depend on correct pre-processing and handling of the data [10].
+We followed the pre-processing approach of Gruver [10] and this process involves the following few steps.
+Input Pre-processing.
+In this phase for time series forecasting with LLMs, we perform two pre-processing steps.
+First, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing.
 For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space.
-Tokenization is equally pivotal, shaping the model’s pattern recognition capabilities.
-Unlike traditional methods like byte-pair encoding (BPE) Hugging Face (2023), which may disrupt numerical coherence, spacing digits ensures individual tokenization, enhancing pattern discernment.
-Moreover, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so a specific percentile aligns to 1, facilitating the model’s exposure to varying digit counts and supporting the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
+Second, tokenization is equally important, shaping the model’s pattern recognition capabilities.
+Unlike traditional methods such as byte-pair encoding (BPE) [13], which can disrupt numerical coherence, we use spacing digits which ensures individual tokenization, enhancing pattern discernment.
+Third, rescaling is employed to efficiently utilize tokens and manage large inputs by adjusting values so that a specific percentile aligns to 1.
+This facilitates the model’s exposure to varying digit counts and supports the generation of larger values, a testament to the nuanced yet critical nature of data preparation in leveraging LLMs for time series analysis.
 2.2 Time Series Forecasting
 In the context of time-series forecasting, the primary goal is to predict the values for the next steps based on observed values from the preceding steps, which is mathematically expressed as:
-Here, represents the -step estimation given the previous -step values .
-denotes the trained parameters from the model , and denotes the prompt or any other information used for the inference.
-This paper predominantly focuses on single-variate time series forecasting.
-Motivated by the interpretability requirements in real-world scenarios, time series can often be decomposed into the trend component, the seasonal component, and the residual component through the addictive model Cleveland et al.
-(1990).
+Here, represent the -step estimation given the previous -step values .
+denotes the trained parameters from the model , and denotes the prompt or any other information used for inference.
+In this paper, we focus predominantly on univariate time series forecasting to investigate the preference and performance of LLMs in univariate time series forecasting under the zero-shot setting.
+Motivated by interpretability requirements in real-world scenarios, time series can often be decomposed into the trend component, the seasonal component, and the residual component through the addictive model [5].
 The trend component captures the hidden long-term changes in the data, such as the linear or exponential pattern.
-The seasonal component captures the repeating variation in the data, and the residual component captures the remaining variation in the data after the removal of trend and seasonal components.
-This decomposition offers a method to quantify the properties of the time series, which is detailed in section 3.2.
+The seasonal component captures the repeating variation in the data, and the residual component captures the remaining variation in the data after removing the trend and seasonal components.
+This decomposition offers a method to quantify the properties of time series, which is detailed in subsection 3.2.
+Datasets.
+In this study, we primarily use Darts [11], a benchmark univariate dataset widely recognized in deep learning research, along with many baseline methods.
+Darts consists of eight real univariate time series datasets, including those with clear patterns, such as the AirPassengerDataset, and irregular datasets, such as the SunspotsDataset.
+Besides, we employ some other commonly used datasets, such as US Births Dataset[9], TSMC-Stock and Turkeypower datasets [10] and ETT [22] in Sections 5.1 and 5.2 to demonstrate the effectiveness of our proposed methods.
+A full description of those datasets can be seen in Appendix A.1.
+Evaluation Metrics.
+In this paper, we evaluate model performance with three metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE).
+These metrics are defined as follows:
+where denotes the true value, represents the predicted value, and is the sample size.
 3 What are LLMs’ Preferences in Time Series Forecasting?
-To explore the preference in the large language model, we first quantify the properties of the input time series to investigate the LLMs’ preferences for time series.
+To explore the preference of LLMs, we first quantify the properties of the input time series to investigate the LLMs’ preferences for time series.
 Then, to further emphasize our findings, we evaluate the importance of different segments of the input sequence by adding Gaussian noise to the original time series.
 3.1 Analyzing Method
-To understand the preferences of the LLMs in time series forecasting, we compare LLMTime using various foundational models like GPT-4 and GPT-3.5-turbo traditional methods on various datasets.
+We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 9 and Table 10.
+It is shown that LLMs perform better within most datasets.
+GPT-4-turbo and Llama-2 perform relatively well on the AirPassengerdataset and the AusBeerdataset with low MAPE.
+Gemini outperforms GPT-3.5-turbo on time series forecasting and outperforms GPT-4-turbo on some datasets but is on par with GPT-4-turbo overall.
+To understand the preferences of the LLMs, we compare our framework using various foundational models, such as GPT-4-turbo and GPT-3.5-turbo, with traditional methods.
 We also design experiments on synthesized datasets to validate our findings and analyze the impact of the multiple periods.
-To quantify the LLMs’ preferences towards time series, following Wang et al.
-(2006), we define the strength of the trend and the seasonality as follows:
+To quantify the LLMs’ preferences towards time series, following [20], we define the strength of the trend and the seasonality as follows:
 where , and denote the trend component, the seasonal component and the residual component respectively.
-The presented indices serve as indicators of the strength of the trend and seasonality, providing a measure ranging up to 1.
+The presented indices indicate the trend’s strength and seasonality, providing a measure ranging up to 1.
 It is easy to find that a higher value indicates a stronger trend or seasonality within the time series.
-To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian to the original time series to create counterfactual examples.
-We initiate by defining a sliding window that constitutes 10% of the total length of the time series.
-Given the variability in the length of time sequences, we first scale the sequence before incorporating the noise into the original sequence.
-This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLMs predominantly focus on.
-3.2 Preferences for the time series Experiment
-In this subsection, we delve into the input sequence preferences for time series forecasting with LLMs.
-Our experiments are conducted on both real datasets and synthetic datasets.
-Throughout this subsection, we use GPT-3.5-turbo-instruct and GPT-4 for time series forecasting and measure model performance through and Mean Absolute Percentage Error (MAPE).
+Throughout this paper, we use the word "higher strength" to represent the comparison of the strengths between different datasets.
+The assessment of strength is not based on a fixed level, as the concepts of "strong" and "weak" vary across different datasets and scenarios.
+To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian noise to the original time series to create counterfactual examples.
+We start by defining a sliding window that constitutes 10% of the total length of the time series, and we set the sliding window to gradually move closer to the output sequence.
+This method allows us to assess the impact of different segments fairly and thereby infer the interpretability of the time series segments that LLMs predominantly focus on.
+3.2 Preferences for Input Sequences
+In this subsection, we investigate the input sequence preferences for time series forecasting with LLMs.
+We conduct experiments on real datasets with GPT-3.5-turbo and GPT-4-turbo, measuring model performance through MAPE.
+To further validate our findings, we also use GPT-3.5-turbo and Gemini-1.0-Pro to forecast multiple-period time series on synthesized datasets.
+| Metrics | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
+| GPT4-MAPE | 1.00000 | 0.987398 | -0.020637 | -0.681440 |
+| GPT3.5-MAPE | 0.987398 | 1.00000 | -0.115087 | -0.669983 |
+| Trend Strength | -0.020637 | -0.115087 | 1.00000 | 0.508980 |
+| Seasonal Strength | -0.681440 | -0.669983 | 0.508980 | 1.00000 |
 3.2.1 Implementation Details
-Real Datasets: We conduct experiments on ten real datasets, as enumerated in Appendix A.2.
-we apply the Seasonal-Trend decomposition using the LOESS (STL) techniqueCleveland et al.
-(1990), to decompose the original time series into trend, seasonal, and residual components.
-Subsequently, we computed the strengths of the trend strength and seasonal strength .
-To further discern the LLMs’ preferences for the specific segments of the input data, we conduct the counterfactual analysis with a systematic permutation to the input time series.
-We initiate by defining a sliding window that constitutes 10% of the total length of the time series and add Gaussian noise into the data within this window data.
-Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence.
-Refer to A.2.1 for detailed information.
-Synthesized Datasets: To further validate our findings, we conduct experiments on synthesized datasets.
-We formulate the dataset through the expression .
-To investigate the influence of the number of periods on model performance, we generated a dataset using the function .
+Real Datasets: We conduct experiments on ten real-world datasets, including both those with clear patterns and those with irregular characteristics.
+The results are shown in Table 6.
+We apply the Seasonal-Trend decomposition using the LOESS (STL) technique [5] to decompose the original time series into trend, seasonal, and residual components.
+Subsequently, we compute the strengths of the trend strength and seasonal strength .
+To further understand the LLMs’ preferences for the specific segments of the input data, we conduct the counterfactual analysis with a systematic permutation to the input time series.
+We first scale the sequence through max-min normalization.
+We then define a sliding window that constitutes 10% of the total length of the time series and add Gaussian noise into the data within this window data.
+Subsequently, the sliding window moves closer to the last known data point.
+Refer to subsubsection A.2.1 for detailed information.
+Synthesized Datasets: To further validate our findings and investigate the influence of the number of periods on model performance, we generate a dataset using the function .
 ranges from 0 to 20 and follows the normal distribution .
-Refer to A.2.1 for detailed information.
+Refer to subsubsection A.2.1 for detailed information.
 3.2.2 Key Findings
-After computing the Pearson correlation coefficients (PCC), we observed a nearly strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series has a higher trend and seasonal strength.
-Notably, GPT-4 achieved a higher PCC compared to GPT-3.5-turbo-instruct.
-It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data.
-Interestingly, there is an increase in the of the output generated by GPT-4, compared to the original test sequences.
-It indicates that GPT-4 tends to forecast time series with high seasonal strength, which may provide insights for further research.
-In the context of multi-period time series, the model performance decreases as the number of periods increases.
-It indicates that LLMs may have trouble recognizing the multiple periods inherent in such datasets, which is common in reality.
-For the counterfactual analysis, as shown in Figure 6, there is a noticeable decrease in values when Gaussian noise is added to the latter segments.
-Our findings reveal that LLMs are more sensitive to the end of input time series data when it acts as a time series predictor.
-4 Why do LLMs forecast well on datasets with higher trend or seasonal strengths?
-Our findings indicate that LLMs exhibit enhanced performance on time series data characterized by pronounced trends or significant seasonal variations.
-This raises the question: Why do LLMs excel in forecasting datasets with marked trend or seasonal patterns?
-To explore this phenomenon, we crafted prompts that necessitate LLMs to acknowledge the temporal span of the dataset.
-This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct trend or seasonal attributes in-context learningvanced pattern recognition capabilities.
-By explicitly prompting LLMs to consider the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, thereby shedding light on the mechanisms that underpin their superior performance in such contexts.
+After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1.In the context of multi-period time series, the model performance worsens as the number of periods increases.
+It indicates that LLMs may have difficulty recognizing the multiple periods inherent in such datasets.
+This potentially stems from their inability to adequately capture long-term periods, which can be supported in Figure 8.
+Besides, for counterfactual analysis, as shown in Figure 2 and Figure 3, there is a noticeable increase in MAPE values when Gaussian noise is added to the latter segments, while the perturbation of the first part of the sequence has little effect on the prediction performance.
+Our findings reveal that LLMs are more sensitive to the end of input time series when forecasting.
+We show our full results in Figure 6 and Figure 7.
+As we move to the right along the x-axis, the closer it gets to the output sequence.
+We defer the details in Appendix A.
+It is also found that the initial part of the sequence has the least impact on the prediction accuracy.
+For the datasets with high seasonal strengths over 85%, such as WoolyDataset, and MonthlymilkDataset, more than 80% of the length of the time series has almost no effect on the model performance.
+4 Why do LLMs Forecast Well on Data with Higher Seasonal Strengths?
+Our findings show that LLMs demonstrate enhanced performance in time series forecasting with strong seasonal strengths.
+This raises the question: Why do LLMs perform well in forecasting datasets with marked seasonal patterns?
+To explore this phenomenon, we craft prompts that require LLMs to recognize the dataset’s temporal pattern.
+This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct seasonal attributes.
+By explicitly prompting LLMs to predict the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, which sheds light on the mechanisms that underpin their superior performance in such contexts.
 4.1 Implementation Details
-In order to explore the phenomenon that LLMs forecast well on datasets with higher trend or seasonal strengths, we design experiments to verify this phenomenon.
-We design prompts to let the LLMs output the predicted value after each sequence prediction.
-The target model of the experiment is GPT-3.5-Turbo (gpt-3.5-turbo-1106), and the main role of our prompt is to perform in-context learning and require output cycles, as detailed in the Appendix A.3.
-We select eight time series data sets such as AirPassengersDataset, count the period value after ten times of prediction, and compare the median of these ten results with the real period.
-The results are shown in Table 3.
-| Dataset Name | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
-| AirPassengersDataset | ||||||
-| AusBeerDataset | ||||||
-| MonthlyMilkDataset | ||||||
-| SunspotsDataset | ||||||
-| WineDataset | ||||||
-| WoolyDataset | ||||||
-| IstanbulTrafficGPT | ||||||
-| GasRateCO2Dataset | ||||||
-| HeartRateDataset | ||||||
-| TurkeyPower |
+To explore the phenomenon that LLMs forecast well on datasets with higher seasonal strengths, we design experiments to verify this phenomenon.
+We tokenize the input sequence and let the LLMs output the period directly.
+We use GPT-3.5-turbo, GPT-4-turbo and Gemini-1.0-Pro to predict the periods.
+We have chosen five datasets with their seasonal strengths exceeding 85%.
+These datasets are readily available with clear seasonal patterns.
+In contrast, determining the specific periods of other irregular datasets is challenging, as they have no specific cycles.
+We record the predicted periods ten times and identify the mode period, which is the most frequently predicted value.
+We then compare the mode of these ten results with the real period.
+The mode is selected as the evaluation metric because, when considering the usage characteristics of LLMs, the output of this number best represents the model’s normal performance.
+We defer our details of the prompt in subsection A.3, and the results are shown in Table 2.
 4.2 Key Findings
-According to the results, we find that large language models can determine the periodicity of a dataset to some extent.
-Although the fluctuation of each time series prediction is large, the prediction on AirPassenger, Sunspots and Wooly datasets is relatively accurate, and the predicted values on other datasets are also close to our real cycle value.
-We speculate that LLMs can predict datasets with high trend or seasonal intensity well because they already possess some knowledge about the scene and content of the dataset.
-5 How to leverage these insights to improve model’s performance?
-Based on these findings, our focus is on how to leverage these findings to further improve model performance.
-Because our ultimate goal is to improve the time series prediction performance of the model.
-Therefore, we retain the original structure of the time series data but enhance the model’s input with prompts infused with additional knowledge.
-These prompts are crafted with seasonal trends and cyclical behaviours in the dataset to endow the model with a richer contextual understanding.
-In another way, we transform time series data into formats that resemble natural language sequences, enabling large language models to leverage their superior inference skills more effectively in the realm of time series prediction.
+According to the results, we find that large language models can mostly determine the periodicity of a dataset.
+The true periods are determined here by the periodogram, which is commonly used to identify the dominant periods [1].
+The multiples of the predicted period also align with the original data cycle.
+Consequently, we consider the prediction of these multiples to be accurate.
+We observe that LLMs generally perform well in predicting the period for most datasets with minimal fluctuations.
+Surprisingly, we discover that in the case of WoolyDataset and AusbeerDataset, which possess relatively short underlying periods, the predicted period is consistently 3 instead of the true period, 4.
+This discrepancy may be attributed to the LLMs’ tendency to focus on cyclic patterns among individual digits rather than considering the entire sequence as a whole, a phenomenon that could also be interpreted as the model’s identification of the underlying cycle.
+We leave a comprehensive analysis of this phenomenon in the future.
+5 How to Leverage These Insights to Improve the Model’s Performance?
+Based on the findings in the previous two sections, our focus is now on how to leverage these findings to further improve model performance.
+In this paper, we propose two approaches to the user input without additional fine-tuning: for the input prompt, we incorporate additional knowledge of the specific trend and seasonal patterns in the dataset, which gives the model a richer understanding of the underlying patterns.
+Regarding the input sequence, we transform the time series data into formats resembling natural language sequences rather than relying on the original tokenization.
+This approach leverages LLMs’ superior capabilities with language sequences.
+Both methods achieve substantially improved model performance.
 5.1 External Knowledge Enhancing Time Series Forecasting
-We introduce a novel method to improve the performance and stability of large language models for time series forecasting.
-The core idea of this part is to use the knowledge obtained from pre-training of large language models to help predict.
-We will give the large language model some basic information about the current dataset such as the name of the dataset in the prompt, and this process does not involve data leakage (it does not tell the large language model about the period or predicted value, etc.) Assuming we define initial prompt as the original time sequence, the extra information defined as , the new prompt can be written as:
+We introduce a novel method to improve the performance of large language models for time series forecasting.
+The core idea of this part is to use the knowledge obtained from the pre-training stage to help predict.
+We provide the large language model with some basic information about the current dataset such as the background of the data collection, and this process does not involve data leakage.
+We incorporate our tests on the data leakage in Appendix A.4.
+It is noted that we do not provide the LLMs with any statistical information such as the periods or trends.
+This approach ensures that the LLMs forecast the time series entirely based on the data and their prior knowledge.
+Let denote the initial prompt representing the original time sequence, and let denote the additional information.
+Consequently, the new prompt can be expressed as: .
 5.1.1 Implementation Details
 We input the dataset’s external knowledge through prompts before the sequence’s input.
-The external knowledge of each dataset is presented in Appendix A.1.
-The external knowledge includes the description of dataset features and scenarios and the dataset’s data properties.
-The results are shown in Table 5.
+The external knowledge of each dataset is presented in subsection A.1.
+The results are shown in Table 8, where LLMTime Prediction refers to the approach described by [10] without any modifications.
 5.1.2 Key Findings
-In External Knowledge Enhancement, GPT-4 generally performs better than GPT-3.5 on MSE, MAE and MAPE, especially on AirPassengers, AusBeer and other datasets.
-Llama-2 significantly outperforms GPT-3.5 and GPT-4 in terms of MSE and MAE metrics on some datasets (e.g., Wooly, ETTh1, ETTm2), indicating that it can capture data features more accurately.
-R-squared values are shown on some datasets such as ETTh1, ETTm2, and Turkey Power, all models can provide relatively accurate predictions with R-squared values close to 1.
-In the LLMTime Prediction, GPT-4 and Llama-2 perform relatively well on AirPassengers and AusBeer datasets with R values close to or above 0.5.
-GPT-3.5 achieves very high MSE on Sunspots and Wine datasets, which may be due to the complexity of the dataset or the poor adaptation of the model for these specific tasks.
-The Turkey Power dataset has a high R value on the time prediction task for all models, indicating a good model fit for this task.
-5.2 Natural Language Paraphrasing of Sequences
-We will be creating experiments related to natural language paraphrasing of sequences.
-This will be based on the fact that LLMs are not affected by the order of magnitude and size of digits Shah et al.
-(2023).
-We have created a method to convert sequences into natural language paraphrases, the main process is using traditional natural language processing techniques to preprocess the series.
-For instance, if we have a discrete time series that represents temperature changes, we can denote it as , where .
-To do this, we extract a natural language description of the sequence based on two data points and a trend from to , "This is a discrete temperature time series.
-The temperature rises from to , and falls from to …".
-The string we get is our natural language paraphrasing sequence.
-The sequence generated by the natural language paraphrasing of the text in is represented by .
-Finally, we paraphrase the sequence from to through the large language model; we let the LLM prediction of get , after Extraction and Processing, get , then the above completed our prediction task.
-We load in the sequence and determine the movement of the sequence at each step.
-Then add "from xx increasing to xx," or "from xx decreasing to xx." Complete the process of modifying from time series to natural language.
-To convert a text sequence into a discrete sequence, we use the Reverse function, which converts the natural language series into a regular expression format through hard code.
+As shown in Table 8, this method achieves improved performance in most scenarios.
+Besides, GPT-4-turbo generally performs better than GPT-3.5-turbo on MSE, MAE, and MAPE, especially on AirPassengers, AusBeer, and other datasets.
+Llama-2 significantly outperforms GPT-3.5-turbo and GPT-4-turbo in terms of MSE and MAE metrics on some datasets (e.g., Wooly, ETTh1, ETTm2), indicating that it can capture data features more accurately.
+Using External Knowledge Enhancing, Gemini outperforms other models on MonthlyMilk, Sunspots, Wooly, and HeartRate Datasets, but performs poorly on other datasets.
+5.2 Natural Language Paraphrasing
+In this subsection, we conduct experiments on the natural language paraphrasing of the input time sequences.
+This strategy capitalizes on the advanced abilities of large language models in handling language sequences.
+It is motivated by the fact that LLMs are insensitive by the order of magnitude and size of digits [18].
+We use natural language to describe the trend between consecutive values.
+For instance, given a time series where , we describe the trend from to as follows: "The value rises from to , and falls from to …".
+The string we get here is our natural language paraphrasing sequence.
+After generating responses based on the string, we extract the values from the text and construct the predicted time series.
+| Model | Dataset | Period | Real | Mode | |||||||||
+| AirPassengersDataset | 24 | 24 | 7 | 24 | 12 | 24 | 11 | 24 | 24 | 24 | 12 | 24 | |
+| WineDataset | 11 | 12 | 24 | 24 | 24 | 20 | 24 | 24 | 24 | 24 | 12 | 24 | |
+| GPT-3.5-turbo | MonthlyMilkDataset | 6 | 9 | 12 | 9 | 12 | 12 | 12 | 12 | 12 | 11 | 12 | 12 |
+| WoolyDataset | 4 | 3 | 4 | 3 | 3 | 4 | 3 | 3 | 6 | 3 | 4 | 3 | |
+| AusBeerDataset | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 4 | 3 | |
+| AirPassengersDataset | 11 | 12 | 12 | 4 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | |
+| WineDataset | 10 | 12 | 24 | 12 | 6 | 12 | 12 | 24 | 12 | 12 | 12 | 12 | |
+| Gemini-Pro-1.0 | MonthlyMilkDataset | 16 | 12 | 12 | 12 | 12 | 39 | 12 | 11 | 12 | 12 | 12 | 12 |
+| WoolyDataset | 5 | 7 | 4 | 5 | 4 | 4 | 4 | 4 | 5 | 6 | 4 | 4 | |
+| AusBeerDataset | 4 | 4 | 4 | 2 | 5 | 5 | 4 | 3 | 5 | 7 | 4 | 4 | |
+| AirPassengersDataset | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | |
+| WineDataset | 7 | 6 | 6 | 6 | 7 | 7 | 6 | 6 | 6 | 6 | 12 | 6 | |
+| GPT-4-turbo | MonthlyMilkDataset | 10 | 12 | 12 | 12 | 12 | 14 | 12 | 12 | 12 | 12 | 12 | 12 |
+| WoolyDataset | 5 | 5 | 7 | 5 | 5 | 5 | 7 | 5 | 5 | 4 | 4 | 5 | |
+| AusBeerDataset | 4 | 4 | 4 | 4 | 6 | 4 | 4 | 4 | 6 | 4 | 4 | 4 | |
 5.2.1 Implementation Details
-We use the LLMTime model with GPT-3.5-Turbo, GPT-4-turbo and Llama-2 as the core to forecast the built-in time series data set, the training set and test set ratio is 4:1.
-We will get the time series prediction results and the original series MSE, MAE, MAPE and R.
-The result is present in Table 4 compared with LLMTime forecasting results.
 5.2.2 Key Findings
-According to the experimental data obtained in the above table, we find that enhancing LLM through natural language paraphrasing has a certain improvement effect on time series analysis and prediction on some datasets.
-From the perspective of model performance comparison, the R value of GPT-3.5-Turbo in Natural Language Paraphrasing and LLMTime Prediction method is negative in some cases, indicating that the model performs poorly.
-GPT-4-Turbo performs better on most datasets compared to other magic methods, especially on Natural Language Paraphrasing methods, MSE, MAE, and MAPE usually decrease, while R improves.
-Llama-2 shows better performance on some datasets, such as lower MSE and MAE and better MAPE and R in the Natural Language Paraphrasing method of the AusBeer dataset.
-6 Related Works
-6.1 Time Series Prediction
-Time series prediction is an important field in the field of artificial intelligence.
-The unified idea of time series prediction is to analyze past and future series trends.
-It is used in finance and business to predict the trend of indexes and stocks, in the field of intelligent manufacturing to predict anomaly detection and power load, and in the field of meteorology, agriculture, and navigation to predict temperature, humidity, and climate Singh and Malhotra (2023); Munir et al.
-(2018); Shen et al.
-(2020); Braei and Wagner (2020).
-We categorize related research within this domain into three categories: traditional time series prediction, time series prediction based on machine learning, and deep learning.
-| Metrics | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
-| GPT4-MAPE | ||||||
-| GPT4-R2 | ||||||
-| GPT3.5-MAPE | ||||||
-| GPT3.5-R2 | ||||||
-| Trend Strength | ||||||
-| Seasonal Strength |
-Two commonly used methods for traditional time series analysis are the ARIMA method Box and Pierce (1970) and the exponential smoothing method Gardner Jr (2006).
-The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions.
-On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations.
-The ARIMA model requires testing the stationarity of data and selecting the right order.
-On the other hand, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower compared to the ARIMA model.
-| Dataset | Period | Real | Median | |||||||||
-| AirPassengersDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
-| AusBeerDataset | 24 | 10 | 6 | 6 | 1 | 3 | 6 | 7 | 3 | 21 | 4 | 6 |
-| GasRateCO2Dataset | 7 | 36 | 24 | 72 | 28 | 34 | 20 | 7 | 16 | 14 | 39 | 20 |
-| MonthlyMilkDataset | 6 | 6 | 25 | 11 | 6 | 2 | 7 | 34 | 8 | 7 | 12 | 7 |
-| SunspotsDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
-| WineDataset | 1 | 11 | 12 | 11 | 12 | 7 | 6 | 8 | 10 | 12 | 12 | 10 |
-| WoolyDataset | 10 | 5 | 6 | 7 | 16 | 11 | 5 | 11 | 3 | 4 | 4 | 5 |
-| HeartRateDataset | 13 | 16 | 6 | 12 | 13 | 11 | 12 | 32 | 32 | 64 | 144 | 12 |
-Time series analysis based on linear regression, decision tree random forest, and other methods can make predictions for future series through the analysis and learning of existing series, as more complex time series prediction methods based on deep learning, RNN, CNN, and LSTM have played a huge role in many applications.
-Especially in the field of sequence prediction, many variants of LSTM have emerged, such as bidirectional LSTM Graves et al.
-(2005), gated recurrent unit Chung et al.
-(2014), long short-term memory network Malhotra et al.
-(2015), and deep recurrent neural network Pascanu et al.
-(2013).
-This method improves prediction accuracy but takes longer to train and is less prone to overfitting than the original LSTM.
-6.2 LLMs for Time Series Prediction
-Given the notable achievements of LLMs across various domains, it is prudent to incorporate time series prediction tasks with LLMs.
-Similar to (Sun et al., 2023), we categorize contemporary research within this domain into two categories: LLM-for-Time Series and Time Series-for-LLM.
-LLM-for-Time Series involves either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data Rasul et al.
-(2023); Garza and Mergenthaler-Canseco (2023); Das et al.
-(2023); Cao et al.
-(2023).
-For instance, (Rasul et al., 2023) aimed to build the foundational models for time series and investigate its scaling behavior.
-Chang et al.
-(2024) proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting.
-Although these studies contribute significantly to understanding foundational models, they demand considerable computing resources and expertise in fine-tuning procedures.
-Moreover, the details of the model may not be disclosed for commercial purposes Garza and Mergenthaler-Canseco (2023), which impedes future research.
-Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
-Therefore, in this paper, we concentrate on the Time Series-foR-LLM, harnessing fixed LLMs for the time-series forecasting task.
+According to the results in Table 3, we find that enhancing LLM through natural language paraphrasing improves time series forecasting on most datasets.
+For instance, GPT-3.5-turbo and GPT-4-turbo perform better on most datasets, especially on Natural Language Paraphrasing methods.
+Gemini outperforms other LLMs on Wooly and Ausbeer datasets but underperforms on others with natural language paraphrasing.
+All these results demonstrate the superior performance of our methods.
 | Models | Datasets | Natural Language Paraphrasing | LLMTime Prediction | ||||
@@
 | Sunspots | 251.61 | 4.27 | 20.42 | 6556.55 | 58.95 | 217.94 | |
-| Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 | |
-| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
-| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
+| (GPT-3.5-turbo-1106) | HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 |
+| Istanbul-Traffic | 224.17 | 3.74 | 8.81 | 335.05 | 6.75 | 11.68 | |
+| ETTh1 | 1.21 | 0.48 | 54.17 | 5.64 | 2.71 | 1.625 | |
+| ETTm2 | 0.81 | 0.36 | 27.33 | 3.46 | 2.17 | 1.178 | |
 | GPT-4-Turbo | AirPassengers | 133.10 | 2.87 | 0.80 | 1286.25 | 28.04 | 6.07 |
@@
 | Sunspots | 194.52 | 5.30 | 16.10 | 3374.70 | 41.87 | 321.11 | |
-| Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 | |
-| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
-| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
-| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1286.25 | 28.04 | 6.07 |
+| (GPT-4-turbo-preview) | HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 |
+| Istanbul-Traffic | 176.91 | 3.88 | 9.67 | 195.33 | 5.53 | 10.03 | |
+| ETTh1 | 1.20 | 0.49 | 47.62 | 4.73 | 1.53 | 3.282 | |
+| ETTm2 | 0.45 | 0.27 | 23.62 | 2.30 | 1.034 | 1.607 | |
+| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1317.9 | 55.49 | 11.18 |
 | AusBeer | 591.75 | 23.25 | 5.41 | 644.82 | 17.88 | 4.08 | |
@@
 | Sunspots | 1483.29 | 33.27 | 17.79 | 4467.67 | 48.95 | 91.79 | |
-| Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 | |
-| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
-| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
-Time Series-for-LLM centres around using existing LLMs and designing mechanisms, such as crafting appropriate prompts or reprogramming inputs, to handle time series data effectively (Gruver et al., 2023; Sun et al., 2023; Jin et al., 2023; Xue and Salim, 2023).
-Sun et al.
-(2023) tokenized the time series and managed to embed those tokens, and Jin et al.
-(2023) reprogrammed the time series data with text prototypes before feeding them to the LLMs.
-These studies illuminated the characteristics of time series data and devised methods to align them with LLMs.
+| (llama-2-13B) | HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 |
+| Istanbul-Traffic | 306.80 | 5.39 | 7.24 | 438.28 | 7.28 | 9.81 | |
+| ETTh1 | 1.47 | 0.87 | 58.34 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 0.84 | 0.41 | 29.86 | 3.31 | 2.07 | 2.153 | |
+| Gemini-Pro-1.0 | AirPassengers | 4474.54 | 31.54 | 7.02 | 6392.21 | 63.57 | 14.03 |
+| AusBeer | 278.45 | 10.05 | 2.29 | 397.78 | 14.36 | 3.27 | |
+| GasRateCO2 | 13.29 | 2.50 | 4.38 | 18.99 | 3.57 | 6.46 | |
+| MonthlyMilk | 440.29 | 11.91 | 1.39 | 628.98 | 17.01 | 1.99 | |
+| Sunspots | 438.29 | 10.47 | 1.21 | 626.03 | 14.94 | 1.73 | |
+| (gemini-1.0-pro) | HeartRate | 40.57 | 4.20 | 4.67 | 57.96 | 6.01 | 6.66 |
+| Istanbul-Traffic | 267.43 | 5.69 | 8.37 | 321.56 | 7.32 | 9.71 | |
+| ETTh1 | 1.17 | 0.74 | 54.86 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
+6 Related Work
+In this section, we review two lines of research that are most relevant to ours.
+6.1 Traditional Time Series Forecasting
+Two commonly used methods for traditional time series analysis are the ARIMA method [2] and the exponential smoothing method [7].
+The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions.
+On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations.
+The ARIMA model requires testing data stationarity and selecting the right order.
+However, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower than the ARIMA model.
+6.2 LLMs for Time Series Forecasting
+The first family of methods involve either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [17, 8, 6, 3].
+For instance, [17] aimed to build the foundational models for time series and investigate its scaling behavior.
+[4] proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting.
+Although these studies contribute significantly to understanding foundational models, they require considerable computing resources and expertise in fine-tuning procedures.
+Moreover, the details of the model may not be disclosed for commercial purposes [8], which impedes future research.
+Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
+In contrast, the second family of methods does not involve model parameter finetuning.
+These methods either create appropriate prompts or reprogramme inputs, to effectively handle time series data [10, 19, 14, 21].
+[19] tokenizes the time series and manages to embed those tokens, and [14] reprogrammed the time series data with text prototypes before feeding them to the LLMs.
+These studies illuminate the characteristics of time series data and devise methods to align them with LLMs.
 However, they lack an analysis of the ability and bias in forecasting time series.
-The most related work to us is Gruver et al.
-(2023), though, it lacks a quantitative analysis of the preference for the time series in LLMs, and it failed to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt.
+The most related work to us is [10], though it lacks a quantitative analysis of the preference for the time series in LLMs, and it fails to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt.
 Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
-7 Conclusions
-Our research has unearthed key preferences of large language models (LLMs) in the domain of time series forecasting, revealing a proclivity for data with distinct trends and seasonal patterns.
+7 Conclusions and Future Work
+In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero shot setting, revealing a proclivity for data with distinct trends and seasonal patterns.
 Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity.
-Our proposed strategies incorporating external knowledge and transforming numbers into natural language have both enhanced LLMs’ predictive accuracy.
-Despite encountering challenges with multi-period time series, our work provides a pivotal foundation for future advancements, offering actionable insights and methods to refine the application of LLMs in time series analysis.
+Besides, our results indicate that LLMs struggle with multi-period time series datasets, as they face difficulty in recognizing the distinct periods within them.
+Our findings also suggest that large language models are more sensitive to the segment of input sequences closer to the last known data than other locations.
+Lastly, experimental results indicate that our proposed strategies of incorporating external knowledge and transforming numerical sequences into natural language formats have yielded substantial improvements in accuracy.
 Limitation
@@
 In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods.
-Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the performance of the model in different data domains.
-These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, which underscores the need to expand future research.
+Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the model’s performance in different data domains.
+These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, underscoring the need to expand future research.
 Appendix A Appendix
-A.1 The External Knowledge incorporated in the Prompt
+A.1 Dataset description and the External Knowledge incorporated in the Prompts
+In this subsection, we briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts.
+Following [10], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets.
+Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods.
+Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
+A.1.1 Darts [11]
 (1)AirPassengersDataset: This is a series of monthly passenger numbers for international flights, where each value is in thousands of passengers for that month.
-(2) AusBeerDataset: This is a quarterly series of beer production, with each value representing the kilolitres of beer produced in that quarter.
+(2)AusBeerDataset: This is a quarterly series of beer production, with each value representing the kiloliters of beer produced in that quarter.
 (3)GasRateCO2Dataset: This time series dataset describes monthly carbon dioxide emissions.
-(4) MonthlyMilkDataset: This is a time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
-(5) SunspotsDataset: This is a dataset that records the number of sunspots in each month, where each data is the number of sunspots in that month.
+(4)MonthlyMilkDataset: This time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
+(5)SunspotsDataset: This dataset records the number of sunspots each month, where each data is the number of sunspots in that month.
 (6)WineDataset: This is a dataset of monthly wine production in Australia, where each figure is the number of wine bottles produced in that month.
@@
 (8)HeartRateDataset: The series contains 1800 uniformly spaced instantaneous heart rate measurements from a single subject.
-(9) Weather: This dataset contains weather-related data such as temperature, precipitation, humidity, etc. It is commonly used in weather forecasting and climate research.
-(10) COVID Deaths: This is a collection of data documenting related deaths during the COVID-19 pandemic, including information on the number of deaths, infection rates, etc., for use in outbreak analysis and health policy development.
-(11) Solar Weekly: This dataset contains weekly data on solar power generation, which can be used to analyze and forecast renewable energy production.
-(12) Tourism Monthly: This dataset records monthly tourism-related data such as the number of tourists, tourism revenue, etc., for tourism trend analysis and prediction.
-(13) Australian Electricity Demand: This deals with electricity demand data in Australia, which can be used for electricity market analysis and grid management.
-(14) Pedestrian Counts: Pedestrian count data are recorded for specific areas, such as city streets, for urban planning and traffic management.
-(15) Traffic Hourly: Provides hourly traffic flow data for traffic forecasting and urban planning.
-(16) Hospital: This dataset may contain hospital-related data such as admission rates, bed occupancy rates, etc., for medical resource planning and public health research.
-(17) Fred MD: It is probably a collection of economic data, such as gross domestic product (GDP), unemployment rate, etc., used for macroeconomic analysis and forecasting.
-(18) Tourism Yearly: Similar to "Tourism Monthly", but provides annual tourism data for longer-term analysis of tourism trends.
-(19) Tourism Quarterly: Provides quarterly tourism data between monthly and annual data for medium-term tourism trend analysis.
-(20) US Births: A dataset that records birth rates in the United States, including information such as the number of births, for use in population research and social policy making.
-(21) NN5 Weekly: This may be a weekly financial or economic dataset, including stock market data, economic indicators, and so on.
-(22) Traffic Weekly: Provides weekly traffic flow data for long-term traffic pattern analysis and planning.
-(23) Saugeenday: This may be a daily degree dataset for a specific region, such as the Saugeen region, and the data involved may include environmental, meteorological, or socioeconomic data.
-(24) CIF 2016: This may be a specific economic or financial time series dataset collected in 2016 for economic forecasting and analysis.
-(25) Bitcoin: This dataset deals with the price and transaction data of Bitcoin and is used for cryptocurrency market analysis and forecasting.
-(26) Sunspot: Data recording sunspot activity for astronomical research and prediction of the solar cycle.
-(27) NN5 Daily: This is similar to NN5 Weekly, but provides daily data that may refer to financial markets or economic indicators.
-A.2 Details of the experiments on the LLMs’ Preferences
+(9)ETTh1: This is a time series dataset containing high-frequency energy data of a certain region in China, which is mainly used for energy load forecasting and related time series analysis research.
+(10)ETTm2: Similar to ETTh1, ETTm2 is also an energy time series dataset, but the data frequency or region covered may be different, which is also used for the analysis and prediction of energy consumption.
+A.1.2 Memorization Datasets [10]
+(11)TSMCStockDataset: This is historical trading data about Taiwan Semiconductor Manufacturing Corporation (TSMC) stock, containing information such as share price, volume, and date, and is commonly used in financial analysis and stock market forecasting research.
+(12)TurkeyPowerDataset: This is a time series dataset on national electricity consumption in Turkey, which records the electricity usage in Turkey over a period of time and is often used to analyze and forecast electricity demand.
+(13)IstanbulTrafficDataset: This dataset offers hourly Traffic Index data for Istanbul from October 2022 to May 2023.
+A.1.3 Monash Datasets [9]
+(14)US Births Dataset: This dataset contains the number of births in the US from 1969 to 1988.
+(15)Saugeen River Flow Dataset: This dataset contains the daily mean flow of the Saugeen River at Walkerton in cubic meters per second from 1915 to 1979.
+A.2 Implementation Details
 In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data.
-We first describe both the real and synthesized datasets we use, and then detail the methods we use to investigate the preferences of LLMs.
+We first describe both the real and synthesized datasets we use and then detail the methods we use to investigate the preferences of LLMs.
 A.2.1 Real Datasets
-We conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTraffic and TurkeyPower, as detailed in A.1.
-We apply the Seasonal-Trend decomposition using the LOESS (STL) technique, to decompose the original time series into trend, seasonal, and residual components.
+We begin by comparing the performance of Large Language Models (LLMs) with traditional time series forecasting methods.
+The results are depicted in Figures Figure 4 and Figure 5, while Tables Table 9 and Table 10 present the computed metrics in tabular format.
+To understand the preferences of LLMs, we conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTrafficDataset and TurkeyPowerDataset.
+We apply the Seasonal-Trend decomposition using the LOESS (STL) technique to decompose the original time series into trend, seasonal, and residual components.
 In those datasets, we obtain the periods through the nature of the data.
-For instance, the number of passengers is collected every month in AirPassengersDataset, and it’s natural to obtain that the period is 12.
-For the datasets without explicit periods, such as the IstanbulTraffic, the period was determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period.
+For instance, the number of passengers is collected monthly in AirPassengersDataset, and it’s natural to obtain that the period is 12.
+For the datasets without explicit periods, such as the IstanbulTrafficDataset, the period is determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period.
 The strengths and the model performance can be seen in Table 6.
 Subsequently, we compute trend strength and seasonal strength to measure all of those components.
-We use and MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 2).
-Notably, GPT-4 achieved a higher absolute PCC compared to GPT-3.5-turbo-instruct.
-It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data.
-Interestingly, upon calculating the strengths of the output sequences, there is an increase in the seasonal component of the output generated by GPT-4, in contrast to the original test sequences.
-This may provide some insights for further research into the characteristics of the LLMs.
-Following this, we conduct the counterfactual analysis with a systematic permutation to the input time series.
-We initiate by defining a sliding window that constitutes 10% of the total length of the time seriesm and add Gaussian noise into the data within this window data.
-Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence.
-The size of the sliding window, , is defined as: , where is the total length of the time series.
-We denote the start point of the window as and the Gaussian noise as , to the data within the sliding window, the modified data point, , is given by:
-where , and is the data point after normalization. The change in Mean Squared Error () is defined as the metric to measure the sensitivity of the LLMs:
-where as , reaches its maximum.
-This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLM predominantly focuses on.
+We use MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 1).
+Notably, GPT-4-turbo achieved a higher absolute PCC compared to GPT-3.5-turbo.
+It may be attributed to human feedback during GPT-4-turbo training, as individuals may be more aware of seasonal and trend data.
+This may provide some insights for further research into the characteristics of the LLMs with time series forecasting.
+We also conduct a counterfactual analysis using a systematic permutation of the input time series.
+We begin by defining a sliding window that determines the length of the periods and add Gaussian noise to the data within this window.
+To reduce costs, we move the window by the period length.
+This method allows us to assess the importance of segments that LLM predominantly focuses on.
 Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
 A.2.2 Synthesized Datasets
-To further validate our findings, we conducted experiments on synthesized datasets.
-We formulate the dataset through the expression , where ranges from 0 to 20, and the noise term follows a normal distribution .
-The coefficients and were uniformly sampled for 10 samples respectively, with and .
-Hence, both trend and seasonal strengths were carefully confined to ,, resulting in more discernible conclusions in the output.
-The chosen frequency, , was consistent across the experiments.
-We uniformly sampled 200 data points.
-Our observations validate that LLMs prefer input sequences with higher strengths and generate time series with higher seasonal strength.
-Furthermore, we observed that the model performance cannot be entirely explained by the linear combination of the strengths, implying that the influence of these strengths is not independent.
 To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components.
@@
 Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution .
-Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure6(a).
+Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure 8.
 This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
-A.3 Detail Prompts
-Request output periodic prompts
-Please output the period size of the predicted sequence as an integer. Output an integer directly.
-In-context learning examples
-For example, the sequence: 0.387952, 8.975192, 5.398713, -6.011139, -9.807413, -0.261663, 9.245404, 5.901009, -6.038525, -9.966693, -0.022652, 9.352823,5.317264, -5.809847, -9.565386, -0.325205, 8.711379, 6.176601, -5.911722, -9.078278.We can get the period number is 5.
-so, try to find out this period:
-A.4 Figures and Tables
+A.3 Detailed Prompts for the Period Prediction
+GPT-3.5-turbo and GPT-4-turbo.
+You are a helpful assistant that specializes in time series analysis.
+The user will provide a sequence.
+The sequence is separated by commas.
+You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence.
+Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence.
+The underlying cycle of the sequence is strictly below 8.
+Please infer the cycle without producing any additional text.
+Again, the sequence is separated by commas.
+Sequence:
+Gemini.
+You are a helpful assistant that specializes in time series analysis.
+The user will provide a sequence.
+The sequence is represented by decimal strings separated by commas.
+You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence.
+Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence.
+The underlying cycle of the sequence is strictly below 8.
+Please infer the cycle without producing any additional text.
+Sequence:
+A.4 Tests on Data Leakage
+Since the detailed datasets used for training GPT-4 and Gemini are not directly accessible, we conducted several experiments to indirectly investigate the data leakage issue.
+First, we performed an "Acknowledge Test" where we asked the LLMs if they recognize the dataset based on its name.
+Next, we conducted a "Series Test" by asking the LLMs to predict the first 20 steps of the datasets.
+Finally, we carried out a "Dataset Detection" test, where we fed the first 20 steps into the LLMs and asked them to identify the dataset.
+All these experiments help us determine whether the LLMs only recognize the name or have detailed knowledge of the datasets.
+As shown in 4, we observe that both GPT and Gemini are generally aware of the content of most datasets.
+However, detailed sequence data is primarily known only for the AirPassengers dataset when using GPT, and this level of detail is not known for other datasets.
+During the experiments on the Dataset Detection, it is found that both GPT and Gemini could identify which dataset is being used based on the first 20 data points of the time series.
+All these results indicate that while they can recognize and identify datasets from limited information, they typically do not possess detailed knowledge of the sequence data for a broader range of datasets.
+| Datasets | Acknowledge Test (GPT) | Acknowledge Test (Gemini) | Series Test (GPT) | Series Test (Gemini) | Dataset Detection (GPT) | Dataset Detection (Gemini) |
+| AirPassengers | Yes | Yes | Yes | No | No | No |
+| AusBeer | No | Yes | No | No | No | No |
+| GasRateCO2 | No | Yes | No | No | No | No |
+| MonthlyMilk | Yes | Yes | No | No | No | No |
+| Sunspots | Yes | Yes | No | No | No | No |
+| Wine | Yes | Yes | No | No | No | No |
+| Wooly | No | No | No | No | No | No |
+| HeartRate | Yes | Yes | No | No | No | No |
+A.5 Computational Cost
+We list the average token length cost associated with external knowledge enhancing and natural language paraphrasing for reference.
+Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy.
+It is noted that Natural Language Paraphrasing is judged one by one through hard coding.
+Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time.
+The results are shown in 5.
+| Datasets | Avg Token Length (ori) | Avg Token Length (EKE) | Avg Token Length (NLP) |
+| AirPassengers | 200 | 224 | 797 |
+| AusBeer | 200 | 220 | 797 |
+| GasRateCO2 | 200 | 211 | 797 |
+| MonthlyMilk | 200 | 218 | 797 |
+| Sunspots | 200 | 217 | 797 |
+| Wine | 200 | 217 | 797 |
+| Wooly | 200 | 216 | 797 |
+| HeartRate | 200 | 214 | 797 |
+| Dataset Name | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
+| AirPassengersDataset | 6.80 | 9.98 | 1.00 | 0.98 |
+| AusBeerDataset | 3.69 | 5.12 | 0.99 | 0.96 |
+| MonthlyMilkDataset | 5.12 | 6.25 | 1.00 | 0.99 |
+| SunspotsDataset | 334.30 | 194.29 | 0.81 | 0.28 |
+| WineDataset | 10.90 | 14.98 | 0.67 | 0.92 |
+| WoolyDataset | 20.41 | 19.26 | 0.96 | 0.82 |
+| IstanbulTrafficGPT | 47.29 | 60.11 | 0.31 | 0.72 |
+| GasRateCO2Dataset | 4.21 | 5.97 | 0.65 | 0.50 |
+| HeartRateDataset | 7.90 | 6.75 | 0.42 | 0.49 |
+| TurkeyPower | 3.36 | 3.52 | 0.90 | 0.88 |
+A.6 Figures and Tables
 The predicted results of AirPassengers, AusBeerDataset, GasRateCO2, HeartRate, Istanbul-Traffic datasets.
 The predicted results of MonthlyMilk, Sunspots, TSMCStock, TurkeyPower, WineDataset, Wooly datasets.
+| Models | Datasets | Natural Language Paraphrasing | LLMTime Prediction | ||||
+| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
+| GPT-3.5-Turbo | AirPassengers | 267.66 | 3.66 | 0.99 | 6244.07 | 61.39 | 14.43 |
+| AusBeer | 598.45 | 5.81 | 1.36 | 841.68 | 23.59 | 5.62 | |
+| GasRateCO2 | 3.16 | 0.46 | 0.85 | 10.88 | 2.66 | 4.73 | |
+| MonthlyMilk | 968.69 | 8.61 | 1.02 | 7507.13 | 66.28 | 112.77 | |
+| Sunspots | 251.61 | 4.27 | 20.42 | 6556.55 | 58.95 | 217.94 | |
+| (GPT-3.5-turbo-1106) | Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 |
+| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
+| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
+| Istanbul-Traffic | 224.17 | 3.74 | 8.81 | 335.05 | 6.75 | 11.68 | |
+| Turkey Power | 24382136.98 | 1843.64 | 4.68 | 3882704.14 | 1315.6 | 3.58 | |
+| ETTh1 | 1.21 | 0.48 | 54.17 | 5.64 | 2.71 | 1.625 | |
+| ETTm2 | 0.81 | 0.36 | 27.33 | 3.46 | 2.17 | 1.178 | |
+| US Births | 926136.72 | 633.37 | 7.67 | 1323052.46 | 904.81 | 9.61 | |
+| Saugeen River Flow | 2870.19 | 18.31 | 19.43 | 4100.27 | 26.16 | 27.76 | |
+| GPT-4-Turbo | AirPassengers | 133.10 | 2.87 | 0.80 | 1286.25 | 28.04 | 6.07 |
+| AusBeer | 661.80 | 7.24 | 1.63 | 513.49 | 18.57 | 4.28 | |
+| GasRateCO2 | 2.28 | 0.41 | 0.75 | 7.27 | 2.32 | 4.18 | |
+| MonthlyMilk | 413.63 | 4.94 | 0.57 | 4442.18 | 50.75 | 172.82 | |
+| Sunspots | 194.52 | 5.30 | 16.10 | 3374.70 | 41.87 | 321.11 | |
+| (GPT-4-turbo-preview) | Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 |
+| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
+| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
+| Istanbul-Traffic | 176.91 | 3.88 | 9.67 | 195.33 | 5.53 | 10.03 | |
+| Turkey Power | 60601807.53 | 3118.43 | 8.03 | 113873.28 | 814.46 | 2.17 | |
+| ETTh1 | 1.20 | 0.49 | 47.62 | 4.73 | 1.53 | 3.282 | |
+| ETTm2 | 0.45 | 0.27 | 23.62 | 2.30 | 1.034 | 1.607 | |
+| US Births | 676264.81 | 501.13 | 4.81 | 966092.59 | 678.55 | 7.12 | |
+| Saugeen River Flow | 3100.57 | 19.18 | 21.86 | 4190.05 | 27.41 | 32.15 | |
+| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1317.9 | 55.49 | 11.18 |
+| AusBeer | 591.75 | 23.25 | 5.41 | 644.82 | 17.88 | 4.08 | |
+| GasRateCO2 | 10.16 | 2.89 | 5.16 | 12.78 | 2.97 | 5.47 | |
+| MonthlyMilk | 851.17 | 84.83 | 9.46 | 3410.20 | 41.40 | 240.25 | |
+| Sunspots | 1483.29 | 33.27 | 17.79 | 4467.67 | 48.95 | 91.79 | |
+| (llama-2-13B) | Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 |
+| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
+| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
+| Istanbul-Traffic | 306.80 | 5.39 | 7.24 | 438.28 | 7.28 | 9.81 | |
+| Turkey Power | 3278744.18 | 2191.34 | 28.76 | 2919773.15 | 1388.10 | 3.70 | |
+| ETTh1 | 1.47 | 0.87 | 58.34 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 0.84 | 0.41 | 29.86 | 3.31 | 2.07 | 2.153 | |
+| US Births | - | - | - | - | - | - | |
+| Saugeen River Flow | - | - | - | - | - | - | |
+| Gemini-Pro-1.0 | AirPassengers | 4474.54 | 31.54 | 7.02 | 6392.21 | 63.57 | 14.03 |
+| AusBeer | 278.45 | 10.05 | 2.29 | 397.78 | 14.36 | 3.27 | |
+| GasRateCO2 | 13.29 | 2.50 | 4.38 | 18.99 | 3.57 | 6.46 | |
+| MonthlyMilk | 440.29 | 11.91 | 1.39 | 628.98 | 17.01 | 1.99 | |
+| Sunspots | 438.29 | 10.47 | 1.21 | 626.03 | 14.94 | 1.73 | |
+| (gemini-1.0-pro) | Wine | 181008.34 | 2557.62 | 10.89 | 258584.78 | 3645.23 | 14.60 |
+| Wooly | 45.44 | 4.47 | 4.93 | 64.92 | 6.39 | 7.04 | |
+| HeartRate | 40.57 | 4.20 | 4.67 | 57.96 | 6.01 | 6.66 | |
+| Istanbul-Traffic | 267.43 | 5.69 | 8.37 | 321.56 | 7.32 | 9.71 | |
+| Turkey Power | 45674.41 | 2973.54 | 11.21 | 103203.37 | 2195.68 | 6.07 | |
+| ETTh1 | 1.17 | 0.74 | 54.86 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
+| US Births | 467743.19 | 440.68 | 4.27 | 687862.05 | 595.51 | 6.28 | |
+| Saugeen River Flow | 2857.23 | 19.34 | 18.80 | 4081.75 | 25.91 | 26.87 | |
 | Models | Dataset | External Knowledge Enhancing | LLMTime Prediction | ||||
@@
 | AirPassengers | 3713.99 | 50.37 | 10.88 | 6244.07 | 61.39 | 14.43 | |
-| AusBeer | 669.01 | 21.82 | 5.12 | 707.43 | 20.44 | 4.80 | |
+| AusBeer | 669.01 | 21.82 | 5.12 | 841.68 | 23.59 | 5.62 | |
 | GasRateCO2 | 16.47 | 3.36 | 5.97 | 10.88 | 2.66 | 4.73 | |
-| HeartRate | 59.83 | 6.44 | 6.75 | 3863.65 | 48.90 | 5.49 | |
 | MonthlyMilk | 4781.26 | 55.45 | 6.25 | 7507.13 | 66.28 | 112.77 | |
-| Sunspots | 7072.42 | 62.61 | 194.29 | 33094824.12 | 4209.81 | 17.69 | |
-| GPT-3.5 | Wine | 24925885.81 | 3548.19 | 14.98 | 820422.05 | 760.03 | 16.51 |
-| (gpt-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 65.95 | 7.05 | 7.47 |
+| Sunspots | 7072.42 | 62.61 | 194.29 | 6556.55 | 58.95 | 217.94 | |
+| GPT-3.5-turbo-1106 | Wine | 24925885.81 | 3548.19 | 14.98 | 30488.60 | 388.28 | 15.83 |
+| (GPT-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 526903.08 | 574.58 | 12.00 |
+| HeartRate | 59.83 | 6.44 | 6.75 | 76.83 | 7.15 | 7.42 | |
 | Istanbul-Traffic | 888.31 | 28.16 | 60.11 | 1321.44 | 48.7 | 7.47 | |
@@
 | ETTm2 | 2.00 | 0.89 | 201.84 | 3.46 | 2.17 | 1.178 | |
-| Wind | 36.05 | 4.99 | inf | 32.07 | 3.76 | 10.41 | |
 | AirPassengers | 1262.24 | 30.54 | 6.80 | 1286.25 | 28.04 | 6.07 | |
@@
 | GasRateCO2 | 6.99 | 2.29 | 4.21 | 7.27 | 2.32 | 4.18 | |
-| HeartRate | 78.99 | 6.96 | 7.90 | 1878.24 | 40.79 | 4.74 | |
 | MonthlyMilk | 2209.33 | 44.02 | 5.12 | 4442.18 | 50.75 | 172.82 | |
-| Sunspots | 4571.92 | 50.24 | 334.30 | 10105667.50 | 2438.74 | 9.85 | |
-| GPT-4 | Wine | 14426570.88 | 2734.41 | 10.90 | 1211754.67 | 1017.44 | 21.85 |
-| (gpt-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 85.22 | 6.89 | 7.77 |
+| Sunspots | 4571.92 | 50.24 | 334.30 | 3374.70 | 41.87 | 321.11 | |
+| GPT-4-turbo-preview | Wine | 14426570.88 | 2734.41 | 10.90 | 22488.17 | 253.08 | 9.98 |
+| (GPT-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 942987.19 | 871.64 | 18.55 |
+| HeartRate | 78.99 | 6.96 | 7.90 | 988.14 | 26.57 | 29.22 | |
 | Istanbul-Traffic | 954.88 | 26.92 | 47.29 | 1291.17 | 32.16 | 6.46 | |
@@
 | ETTm2 | 1.18 | 0.79 | 291.67 | 2.30 | 1.034 | 1.607 | |
-| Wind | 1.82 | 0.81 | 82.50 | 28.83 | 6.32 | 2.14 | |
 | AirPassengers | 3713.99 | 50.37 | 10.88 | 1286.25 | 28.04 | 6.07 | |
@@
 | GasRateCO2 | 11.38 | 3.04 | 5.49 | 12.78 | 2.97 | 5.47 | |
-| HeartRate | 112.17 | 7.86 | 8.93 | 3317.83 | 52.71 | 6.15 | |
 | MonthlyMilk | 4722.32 | 60.36 | 7.05 | 3410.20 | 41.40 | 240.25 | |
-| Sunspots | 4000.19 | 46.45 | 138.69 | 11788774.81 | 2770.67 | 10.90 | |
-| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 636637.46 | 742.67 | 15.54 |
-| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 64.92 | 6.39 | 7.04 |
+| Sunspots | 4000.19 | 46.45 | 138.69 | 4467.67 | 48.95 | 91.79 | |
+| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 951194.94 | 240.08 | 9.45 |
+| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 675062.52 | 736.04 | 15.83 |
+| HeartRate | 112.17 | 7.86 | 8.93 | 75.58 | 7.11 | 7.94 | |
 | Istanbul-Traffic | 979.15 | 26.70 | 45.57 | 1531.37 | 34.74 | 7.42 | |
@@
 | ETTm2 | 3.08 | 1.47 | 810.56 | 3.31 | 2.07 | 2.153 | |
-| Wind | 1.78 | 0.79 | 116.31 | 36.78 | 4.02 | 12.38 | |
+| AirPassengers | 5237.85 | 51.92 | 11.08 | 6392.21 | 63.57 | 14.03 | |
+| AusBeer | 325.45 | 10.84 | 1.86 | 397.78 | 14.36 | 3.27 | |
+| GasRateCO2 | 15.54 | 3.23 | 4.43 | 18.99 | 3.57 | 6.46 | |
+| MonthlyMilk | 491.26 | 15.18 | 1.13 | 628.98 | 17.01 | 1.99 | |
+| Sunspots | 491.64 | 11.15 | 1.27 | 626.03 | 14.94 | 1.73 | |
+| Gemini-1.0-pro | Wine | 210818.24 | 3230.41 | 8.35 | 258584.78 | 3645.23 | 14.60 |
+| (gemini-1.0-pro) | Wooly | 51.04 | 5.70 | 7.93 | 64.92 | 6.39 | 7.04 |
+| HeartRate | 47.45 | 4.83 | 4.67 | 57.96 | 6.01 | 6.66 | |
+| Istanbul-Traffic | 1253.74 | 28.25 | 5.42 | 1531.37 | 34.74 | 7.42 | |
+| TSMC-Stock | 153.73 | 5.02 | 1.05 | 188.18 | 6.67 | 1.65 | |
+| Turkey Power | 83522.95 | 1812.31 | 5.51 | 103203.37 | 2195.68 | 6.07 | |
+| ETTh1 | 2.92 | 1.45 | 2.88 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 2.00 | 1.74 | 1.22 | 3.31 | 2.07 | 2.153 | |
 | Dataset | Method | MSE | MAE | MAPE |
@@
 | Naive Moving Average | 6032.80 | 62.87 | 14.18 | |
-| LLMTime with GPT-3.5-Turbo | 267.66 | 3.66 | 0.99 | |
-| LLMTime with GPT-4-Turbo | 133.10 | 2.87 | 0.80 | |
+| N-Beats | 3994.55 | 54.95 | 12.81 | |
+| DeepAR | 184222.64 | 421.99 | 98.42 | |
+| Prophet | 7345.31 | 43.87 | 8.62 | |
+| LLMTime with GPT-3.5-Turbo | 6244.07 | 61.39 | 14.43 | |
+| LLMTime with GPT-4-Turbo | 1317.9 | 55.49 | 11.18 | |
+| LLMTime with Gemini-1.0-pro | 6392.21 | 63.57 | 14.03 | |
+| LLMtime with Llama-2 | 1286.25 | 28.04 | 6.07 | |
 | AusBeer | Exponential Smoothing | 703.26 | 22.80 | 5.44 |
@@
 | Naive Moving Average | 10828.02 | 96.35 | 23.39 | |
-| LLMTime with GPT-3.5-Turbo | 598.45 | 5.81 | 1.36 | |
-| LLMTime with GPT-4-Turbo | 661.80 | 7.24 | 1.63 | |
+| N-Beats | 250.61 | 14.42 | 3.53 | |
+| DeepAR | 16197.17 | 40.23 | 9.89 | |
+| Prophet | 6323.89 | 28.76 | 6.92 | |
+| LLMTime with GPT-3.5-Turbo | 841.68 | 23.59 | 5.62 | |
+| LLMTime with GPT-4-Turbo | 513.49 | 18.57 | 4.28 | |
+| LLMTime with Gemini-1.0-pro | 397.78 | 14.36 | 3.27 | |
+| LLMtime with Llama-2 | 644.82 | 17.88 | 4.08 | |
 | MonthlyMilk | Exponential Smoothing | 564.94 | 20.23 | 2.41 |
@@
 | Naive Moving Average | 4870.40 | 56.00 | 6.31 | |
-| LLMTime with GPT-3.5-Turbo | 968.69 | 8.61 | 1.02 | |
-| LLMTime with GPT-4-Turbo | 413.63 | 4.94 | 0.57 | |
+| N-Beats | 3140.89 | 51.57 | 6.07 | |
+| DeepAR | 728289.50 | 851.30 | 99.22 | |
+| Prophet | 663.41 | 25.76 | 2.92 | |
+| LLMTime with GPT-3.5-Turbo | 7507.13 | 66.28 | 112.77 | |
+| LLMTime with GPT-4-Turbo | 4442.18 | 50.75 | 172.82 | |
+| LLMTime with Gemini-1.0-pro | 628.98 | 17.01 | 1.99 | |
+| LLMtime with Llama-2 | 3410.20 | 41.40 | 240.25 | |
 | Sunspots | Moving Average | 326750.49 | 499.78 | 3129.63 |
@@
 | Naive Moving Average | 4440.63 | 56.78 | 688.58 | |
-| LLMTime with GPT-3.5-Turbo | 251.61 | 4.27 | 20.42 | |
-| LLMTime with GPT-4-Turbo | 194.52 | 5.30 | 16.10 |
-| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
-| AirPassengersDataset | ||||||
-| AusBeerDataset | ||||||
-| GasRateCO2Dataset | ||||||
-| MonthlyMilkDataset | ||||||
-| SunspotsDataset | ||||||
-| WineDataset | ||||||
-| WoolyDataset | ||||||
-| HeartRateDataset | ||||||
-| IstanbulTraffic | ||||||
-| TurkeyPower |
-| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
-| AirPassengersDataset | ||||||
-| AusBeerDataset | ||||||
-| GasRateCO2Dataset | ||||||
-| MonthlyMilkDataset | ||||||
-| SunspotsDataset | ||||||
-| WineDataset | ||||||
-| WoolyDataset | ||||||
-| HeartRateDataset | ||||||
-| IstanbulTraffic | ||||||
-| TurkeyPower |
+| N-Beats | 4877.59 | 56.58 | 105.55 | |
+| DeepAR | 3421.02 | 48.93 | 132.76 | |
+| Prophet | 6303.57 | 76.83 | 67.97 | |
+| LLMTime with GPT-3.5-Turbo | 6556.55 | 58.95 | 217.94 | |
+| LLMTime with GPT-4-Turbo | 3374.70 | 41.87 | 321.11 | |
+| LLMTime with Gemini-1.0-pro | 626.03 | 14.94 | 1.73 | |
+| LLMtime with Llama-2 | 4467.67 | 48.95 | 91.79 |
+| Dataset | Method | MSE | MAE | MAPE |
+| WineDataset | Exponential Smoothing | 23709576.52 | 3370.78 | 14.23 |
+| SARIMA | 1150166.94 | 966.57 | 20.76 | |
+| Cyclical Regression | 7873785.27 | 2148.24 | 8.52 | |
+| AutoARIMA | 698661.90 | 646.03 | 14.07 | |
+| FFT | 1031170.45 | 867.83 | 18.60 | |
+| StatsForecastAutoARIMA | 20040877.37 | 2853.17 | 12.05 | |
+| Naive Mean | 11557786.19 | 2200.04 | 8.80 | |
+| Naive Seasonal | 879447.22 | 724.23 | 15.52 | |
+| Naive Drift | 9609576.04 | 1833.38 | 7.36 | |
+| Naive Moving Average | 9070696.99 | 1719.17 | 6.90 | |
+| N-Beats | 5418377.00 | 1887.30 | 7.68 | |
+| DeepAR | 715027008.00 | 26236.14 | 89.91 | |
+| Prophet | 4846922.27 | 2201.57 | 8.27 | |
+| LLMTime with GPT-3.5-Turbo | 30488.60 | 388.28 | 15.83 | |
+| LLMTime with GPT-4-Turbo | 22488.17 | 253.08 | 9.98 | |
+| LLMTime with Gemini-1.0-pro | 258584.78 | 3645.23 | 14.60 | |
+| LLMtime with Llama-2 | 951194.94 | 240.08 | 9.45 | |
+| WoolyDataset | Exponential Smoothing | 24925885.81 | 3548.19 | 14.98 |
+| SARIMA | 812352.21 | 759.07 | 16.37 | |
+| Cyclical Regression | 1032574.82 | 962.72 | 22.14 | |
+| AutoARIMA | 838852.91 | 786.25 | 16.84 | |
+| FFT | 1012255.35 | 945.20 | 20.80 | |
+| StatsForecastAutoARIMA | 917617.19 | 858.57 | 18.91 | |
+| Naive Mean | 816762.31 | 764.73 | 16.12 | |
+| Naive Seasonal | 1051110.81 | 982.25 | 22.19 | |
+| Naive Drift | 812352.21 | 759.07 | 16.37 | |
+| Naive Moving Average | 1032574.82 | 962.72 | 22.14 | |
+| N-Beats | 653104.31 | 743.54 | 15.96 | |
+| DeepAR | 243831.14 | 4897.85 | 94.89 | |
+| Prophet | 365241.98 | 891.70 | 34.65 | |
+| LLMTime with GPT-3.5-Turbo | 526903.08 | 574.58 | 12.00 | |
+| LLMTime with GPT-4-Turbo | 942987.19 | 871.64 | 18.55 | |
+| LLMTime with Gemini-1.0-pro | 64.92 | 6.39 | 7.04 | |
+| LLMtime with Llama-2 | 675062.52 | 736.04 | 15.83 | |
+| HeartRateDataset | Exponential Smoothing | 11.16 | 1.38 | 1.49 |
+| SARIMA | 12.98 | 1.34 | 1.61 | |
+| Cyclical Regression | 13.58 | 1.31 | 1.20 | |
+| AutoARIMA | 13.26 | 1.25 | 1.39 | |
+| FFT | 13.95 | 1.16 | 1.34 | |
+| StatsForecastAutoARIMA | 10.53 | 1.27 | 1.39 | |
+| Naive Mean | 12.02 | 1.27 | 1.26 | |
+| Naive Seasonal | 10.55 | 1.32 | 1.31 | |
+| Naive Drift | 10.60 | 1.15 | 1.30 | |
+| Naive Moving Average | 12.13 | 1.27 | 1.34 | |
+| N-Beats | 72.11 | 7.10 | 7.40 | |
+| DeepAR | 286.82 | 15.67 | 16.36 | |
+| Prophet | 88.93 | 10.97 | 6.54 | |
+| LLMTime with GPT-3.5-Turbo | 76.83 | 7.15 | 7.42 | |
+| LLMTime with GPT-4-Turbo | 988.14 | 26.57 | 29.22 | |
+| LLMTime with Gemini-1.0-pro | 57.96 | 6.01 | 6.66 | |
+| LLMtime with Llama-2 | 75.58 | 7.11 | 7.94 | |
+| Weather | Exponential Smoothing | 1684.38 | 31.60 | 6.79 |
+| SARIMA | 1943.81 | 33.33 | 7.09 | |
+| Cyclical Regression | 1700.73 | 30.77 | 7.15 | |
+| AutoARIMA | 7315.10 | 57.44 | 11.70 | |
+| FFT | 2752.02 | 38.90 | 8.87 | |
+| StatsForecastAutoARIMA | 2479.55 | 38.06 | 8.16 | |
+| Naive Mean | 39879.84 | 168.27 | 36.44 | |
+| Naive Seasonal | 5057.47 | 52.81 | 11.89 | |
+| Naive Drift | 5466.23 | 60.58 | 14.70 | |
+| Naive Moving Average | 5057.47 | 52.81 | 11.89 | |
+| N-Beats | 4532.84 | 39.21 | 23.49 | |
+| DeepAR | 6325.75 | 35.97 | 16.59 | |
+| Prophet | 3768.15 | 29.36 | 24.01 | |
+| LLMTime with GPT-3.5-Turbo | 224.54 | 3.07 | 0.83 | |
+| LLMTime with GPT-4-Turbo | 111.65 | 2.40 | 0.64 | |
+| LLMTime with Gemini-1.0-pro | 176.32 | 3.72 | 0.75 | |
+| LLMtime with Llama-2 | 215.39 | 4.07 | 1.31 |

### v4

@@
 Recent works treat large language models as zero-shot time series reasoners without further fine-tuning, which achieves remarkable performance.
-However, there are some unexplored research problems when applying LLMs for time series forecasting under the zero-shot setting.
+However, some unexplored research problems exist when applying LLMs for time series forecasting under the zero-shot setting.
 For instance, the LLMs’ preferences for the input time series are less understood.
@@
 This observation can be explained by the ability of LLMs to recognize the underlying period within datasets, which is supported by our experiments.
-In addition, the input strategy is investigated and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series.
-Overall, our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
+In addition, the input strategy is investigated, and it is found that incorporating external knowledge and adopting natural language paraphrases substantially improve the predictive performance of LLMs for time series.
+Our study contributes insight into LLMs’ advantages and limitations in time series forecasting under different conditions.
 1 Introduction
-Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [16, 15, 12].
+Recently, large language models (LLMs) have been widely used and have achieved promising performance across various domains, such as health management, customer analysis, and text feature mining [1, 2, 3, 4].
 Time series forecasting requires extrapolation from sequential observations.
-Language models are designed to discern intricate concepts within temporally correlated sequences, and intuitively appear well-suited for this task.
-Hence, there exist some preliminary studies that apply LLMs to time series forecasting tasks [10, 17, 19].
-However, currently the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined.
+Language models are designed to discern intricate concepts within temporally correlated sequences and intuitively appear well-suited for this task.
+Hence, some preliminary studies apply LLMs to time series forecasting tasks [5, 6, 7].
+However, the application of LLMs for time series forecasting is still in its early stage, and the boundaries of this research area are not yet well defined.
 There are many unexplored problems in this field.
@@
 This includes the effectiveness gap for LLMs in predicting data with seasonal and trending patterns versus data without such patterns.
-To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero shot prompting setting.
-Through experiments on both real and synthesized datasets, we find that LLMs perform better in time series with higher trend or seasonal strengths.
-Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture the distinct periods within those datasets.
-To further discern the LLMs’ preferences for the specific segments of the input data, we design counterfactual experiments involving systematic permutations of input sequences.
+To fill this research gap, in this paper, we focus on LLMs’ preferences for the input time series in time series forecasting under the zero-shot prompting setting.
+Experiments on both real and synthesized datasets show that LLMs perform better in time series with higher trend or seasonal strengths.
+Our observations also reveal that LLMs perform worse when there are multiple periods within datasets, which may be attributed to the fact that LLMs cannot capture distinct periods within those datasets.
+To further discern the LLMs’ preferences for the specific input data segments, we design counterfactual experiments involving systematic permutations of input sequences.
 The findings suggest that LLMs are particularly sensitive to the segment of input sequences closest to the target output.
-Based on the above findings, we want to further explore why LLMs forecast well on datasets with higher seasonal strengths.
+Based on the above findings, we want to explore why LLMs forecast well on datasets with higher seasonal strengths.
 To this end, we require LLMs to tell the period of the datasets through multiple runs.
 We find that LLMs can mostly recognize the underlying period of a dataset.
-This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well, since they can obtain the seasonal pattern inside the datasets.
+This can explain the findings of why large language models can forecast time series with high trends or seasonal intensities well since they can obtain the seasonal pattern inside the datasets.
 In light of the above-mentioned findings, we are interested in how to leverage these insights to further improve model performance.
@@
 Input Pre-processing.
-In this phase for time series forecasting with LLMs, we perform two pre-processing steps.
+In this phase of time series forecasting with LLMs, we perform two pre-processing steps.
 First, numerical values are transformed into strings, a crucial step that significantly influences the model’s comprehension and data processing.
-For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to "1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0", introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space.
+For instance, a series like 0.123, 1.23, 12.3, 123.0 is reformatted to ”1 2, 1 2 3, 1 2 3 0, 1 2 3 0 0”, introducing spaces between digits and commas to delineate time steps, while decimal points are omitted to save token space.
 Second, tokenization is equally important, shaping the model’s pattern recognition capabilities.
@@
 Darts consists of eight real univariate time series datasets, including those with clear patterns, such as the AirPassengerDataset, and irregular datasets, such as the SunspotsDataset.
-Besides, we employ some other commonly used datasets, such as US Births Dataset[9], TSMC-Stock and Turkeypower datasets [10] and ETT [22] in Sections 5.1 and 5.2 to demonstrate the effectiveness of our proposed methods.
-A full description of those datasets can be seen in Appendix A.1.
+Besides, we employ some other commonly used datasets, such as US Births Dataset[11], TSMC-Stock and Turkeypower datasets [5] and ETT [12] in Sections 5.2 and 5.3 to demonstrate the effectiveness of our proposed methods.
+A full description of those datasets can be seen in Section 5.1.
 Evaluation Metrics.
@@
 3.1 Analyzing Method
-We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 9 and Table 10.
+We first compare the performance between LLMs and traditional time series forecasting methods, as shown in Table 2.
 It is shown that LLMs perform better within most datasets.
@@
 It is easy to find that a higher value indicates a stronger trend or seasonality within the time series.
-Throughout this paper, we use the word "higher strength" to represent the comparison of the strengths between different datasets.
-The assessment of strength is not based on a fixed level, as the concepts of "strong" and "weak" vary across different datasets and scenarios.
+Throughout this paper, we use the word ”higher strength” to represent the comparison of the strengths between different datasets.
+The assessment of strength is not based on a fixed level, as the concepts of ”strong” and ”weak” vary across different datasets and scenarios.
 To further discern the LLMs’ preferences for the specific segments of the input data, we add Gaussian noise to the original time series to create counterfactual examples.
@@
 | Metrics | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
-| GPT4-MAPE | 1.00000 | 0.987398 | -0.020637 | -0.681440 |
-| GPT3.5-MAPE | 0.987398 | 1.00000 | -0.115087 | -0.669983 |
-| Trend Strength | -0.020637 | -0.115087 | 1.00000 | 0.508980 |
-| Seasonal Strength | -0.681440 | -0.669983 | 0.508980 | 1.00000 |
+| GPT4-MAPE | ||||
+| GPT3.5-MAPE | ||||
+| Trend Strength | ||||
+| Seasonal Strength |
 3.2.1 Implementation Details
@@
 Subsequently, the sliding window moves closer to the last known data point.
-Refer to subsubsection A.2.1 for detailed information.
 Synthesized Datasets: To further validate our findings and investigate the influence of the number of periods on model performance, we generate a dataset using the function .
 ranges from 0 to 20 and follows the normal distribution .
-Refer to subsubsection A.2.1 for detailed information.
 3.2.2 Key Findings
-After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1.In the context of multi-period time series, the model performance worsens as the number of periods increases.
+After computing the Pearson correlation coefficients (PCC), we observe a nearly strong correlation between the strengths and model performance, showing that LLMs perform better when the input time series has a higher trend and seasonal strength, which is shown in Table 1.
+In the context of multi-period time series, the model performance worsens as the number of periods increases.
 It indicates that LLMs may have difficulty recognizing the multiple periods inherent in such datasets.
-This potentially stems from their inability to adequately capture long-term periods, which can be supported in Figure 8.
 Besides, for counterfactual analysis, as shown in Figure 2 and Figure 3, there is a noticeable increase in MAPE values when Gaussian noise is added to the latter segments, while the perturbation of the first part of the sequence has little effect on the prediction performance.
@@
 As we move to the right along the x-axis, the closer it gets to the output sequence.
-We defer the details in Appendix A.
 It is also found that the initial part of the sequence has the least impact on the prediction accuracy.
 For the datasets with high seasonal strengths over 85%, such as WoolyDataset, and MonthlymilkDataset, more than 80% of the length of the time series has almost no effect on the model performance.
-4 Why do LLMs Forecast Well on Data with Higher Seasonal Strengths?
+4 Why do LLMs Forecast Well on Data with Higher
+Seasonal Strengths?
 Our findings show that LLMs demonstrate enhanced performance in time series forecasting with strong seasonal strengths.
@@
 To explore this phenomenon, we craft prompts that require LLMs to recognize the dataset’s temporal pattern.
-This approach is grounded in the hypothesis that LLMs’ proficiency in handling datasets with distinct seasonal attributes.
+This approach is grounded in the hypothesis that LLMs are proficient in handling datasets with distinct seasonal attributes.
 By explicitly prompting LLMs to predict the dataset’s period, we aim to leverage their inherent ability to discern and extrapolate from complex patterns, which sheds light on the mechanisms that underpin their superior performance in such contexts.
+| AirPassengers | AusBeer | ||
+| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
+| Exponential Smoothing | 2007.67 / 37.91 / 8.10 | Exponential Smoothing | 703.26 / 22.80 / 5.44 |
+| SARIMA | 2320.47 / 39.80 / 8.46 | SARIMA | 475.53 / 19.07 / 4.49 |
+| Cyclical Regression | 2028.37 / 36.70 / 8.52 | Cyclical Regression | 989.31 / 26.29 / 6.13 |
+| AutoARIMA | 8702.09 / 68.52 / 13.98 | AutoARIMA | 550.05 / 18.84 / 4.41 |
+| FFT | 3274.46 / 46.38 / 10.59 | FFT | 7682.56 / 73.74 / 17.44 |
+| StatsForecastAutoARIMA | 2952.52 / 45.41 / 9.71 | StatsForecastAutoARIMA | 559.46 / 20.56 / 4.86 |
+| Naive Mean | 47703.65 / 204.25 / 44.61 | Naive Mean | 1885.72 / 30.66 / 6.68 |
+| Naive Seasonal | 6032.80 / 62.87 / 14.18 | Naive Seasonal | 10828.02 / 96.35 / 23.39 |
+| Naive Drift | 6505.79 / 72.21 / 17.50 | Naive Drift | 18507.61 / 128.23 / 30.91 |
+| Naive Moving Average | 6032.80 / 62.87 / 14.18 | Naive Moving Average | 10828.02 / 96.35 / 23.39 |
+| N-Beats | 3994.55 / 54.95 / 12.81 | N-Beats | 250.61 / 14.42 / 3.53 |
+| DeepAR | 184222.64 / 421.99 / 98.42 | DeepAR | 16197.17 / 40.23 / 9.89 |
+| Prophet | 7345.31 / 43.87 / 8.62 | Prophet | 6323.89 / 28.76 / 6.92 |
+| LLMTime with GPT-3.5-Turbo | 6244.07 / 61.39 / 14.43 | LLMTime with GPT-3.5-Turbo | 841.68 / 23.59 / 5.62 |
+| LLMTime with GPT-4-Turbo | 1317.9 / 55.49 / 11.18 | LLMTime with GPT-4-Turbo | 513.49 / 18.57 / 4.28 |
+| LLMTime with Gemini-1.0-pro | 6392.21 / 63.57 / 14.03 | LLMTime with Gemini-1.0-pro | 397.78 / 14.36 / 3.27 |
+| LLMtime with Llama-2 | 1286.25 / 28.04 / 6.07 | LLMtime with Llama-2 | 644.82 / 17.88 / 4.08 |
+| MonthlyMilk | Sunspots | ||
+| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
+| Exponential Smoothing | 564.94 / 20.23 / 2.41 | Exponential Smoothing | 326750.49 / 499.78 / 3129.63 |
+| SARIMA | 1289.76 / 32.78 / 3.87 | SARIMA | 2902.72 / 45.75 / 466.99 |
+| Cyclical Regression | 3631.53 / 56.15 / 6.60 | Cyclical Regression | 3917.76 / 47.84 / 274.31 |
+| AutoARIMA | 2682.67 / 42.82 / 5.20 | AutoARIMA | 4695.67 / 58.47 / 709.23 |
+| FFT | 3453.96 / 45.62 / 5.48 | FFT | 3784.56 / 49.81 / 150.32 |
+| StatsForecastAutoARIMA | 186.14 / 10.64 / 1.28 | StatsForecastAutoARIMA | 8406.55 / 72.99 / 95.18 |
+| Naive Mean | 19893.07 / 127.33 / 14.46 | Naive Mean | 4120.40 / 49.84 / 267.22 |
+| Naive Seasonal | 4870.40 / 56.00 / 6.31 | Naive Seasonal | 4440.63 / 56.78 / 688.58 |
+| Naive Drift | 3998.11 / 56.06 / 6.52 | Naive Drift | 5032.77 / 60.40 / 724.88 |
+| Naive Moving Average | 4870.40 / 56.00 / 6.31 | Naive Moving Average | 4440.63 / 56.78 / 688.58 |
+| N-Beats | 3140.89 / 51.57 / 6.07 | N-Beats | 4877.59 / 56.58 / 105.55 |
+| DeepAR | 728289.50 / 851.30 / 99.22 | DeepAR | 3421.02 / 48.93 / 132.76 |
+| Prophet | 663.41 / 25.76 / 2.92 | Prophet | 6303.57 / 76.83 / 67.97 |
+| LLMTime with GPT-3.5-Turbo | 7507.13 / 66.28 / 112.77 | LLMTime with GPT-3.5-Turbo | 6556.55 / 58.95 / 217.94 |
+| LLMTime with GPT-4-Turbo | 4442.18 / 50.75 / 172.82 | LLMTime with GPT-4-Turbo | 3374.70 / 41.87 / 321.11 |
+| LLMTime with Gemini-1.0-pro | 628.98 / 17.01 / 1.99 | LLMTime with Gemini-1.0-pro | 626.03 / 14.94 / 1.73 |
+| LLMtime with Llama-2 | 3410.20 / 41.40 / 240.25 | LLMtime with Llama-2 | 4467.67 / 48.95 / 91.79 |
+| WineDataset | WoolyDataset | ||
+| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
+| Exponential Smoothing | 23709576.52 / 3370.78 / 14.23 | Exponential Smoothing | 24925885.81 / 3548.19 / 14.98 |
+| SARIMA | 1150166.94 / 966.57 / 20.76 | SARIMA | 812352.21 / 759.07 / 16.37 |
+| Cyclical Regression | 7873785.27 / 2148.24 / 8.52 | Cyclical Regression | 1032574.82 / 962.72 / 22.14 |
+| AutoARIMA | 698661.90 / 646.03 / 14.07 | AutoARIMA | 838852.91 / 786.25 / 16.84 |
+| FFT | 1031170.45 / 867.83 / 18.60 | FFT | 1012255.35 / 945.20 / 20.80 |
+| StatsForecastAutoARIMA | 20040877.37 / 2853.17 / 12.05 | StatsForecastAutoARIMA | 917617.19 / 858.57 / 18.91 |
+| Naive Mean | 11557786.19 / 2200.04 / 8.80 | Naive Mean | 816762.31 / 764.73 / 16.12 |
+| Naive Seasonal | 879447.22 / 724.23 / 15.52 | Naive Seasonal | 1051110.81 / 982.25 / 22.19 |
+| Naive Drift | 9609576.04 / 1833.38 / 7.36 | Naive Drift | 812352.21 / 759.07 / 16.37 |
+| Naive Moving Average | 9070696.99 / 1719.17 / 6.90 | Naive Moving Average | 1032574.82 / 962.72 / 22.14 |
+| N-Beats | 5418377.00 / 1887.30 / 7.68 | N-Beats | 653104.31 / 743.54 / 15.96 |
+| DeepAR | 715027008.00 / 26236.14 / 89.91 | DeepAR | 243831.14 / 4897.85 / 94.89 |
+| Prophet | 4846922.27 / 2201.57 / 8.27 | Prophet | 365241.98 / 891.70 / 34.65 |
+| LLMTime with GPT-3.5-Turbo | 30488.60 / 388.28 / 15.83 | LLMTime with GPT-3.5-Turbo | 526903.08 / 574.58 / 12.00 |
+| LLMTime with GPT-4-Turbo | 22488.17 / 253.08 / 9.98 | LLMTime with GPT-4-Turbo | 942987.19 / 871.64 / 18.55 |
+| LLMTime with Gemini-1.0-pro | 258584.78 / 3645.23 / 14.60 | LLMTime with Gemini-1.0-pro | 64.92 / 6.39 / 7.04 |
+| LLMtime with Llama-2 | 951194.94 / 240.08 / 9.45 | LLMtime with Llama-2 | 675062.52 / 736.04 / 15.83 |
+| HeartRateDataset | Weather | ||
+| Method | MSE / MAE / MAPE | Method | MSE / MAE / MAPE |
+| Exponential Smoothing | 11.16 / 1.38 / 1.49 | Exponential Smoothing | 1684.38 / 31.60 / 6.79 |
+| SARIMA | 12.98 / 1.34 / 1.61 | SARIMA | 1943.81 / 33.33 / 7.09 |
+| Cyclical Regression | 13.58 / 1.31 / 1.20 | Cyclical Regression | 1700.73 / 30.77 / 7.15 |
+| AutoARIMA | 13.26 / 1.25 / 1.39 | AutoARIMA | 7315.10 / 57.44 / 11.70 |
+| FFT | 13.95 / 1.16 / 1.34 | FFT | 2752.02 / 38.90 / 8.87 |
+| StatsForecastAutoARIMA | 10.53 / 1.27 / 1.39 | StatsForecastAutoARIMA | 2479.55 / 38.06 / 8.16 |
+| Naive Mean | 12.02 / 1.27 / 1.26 | Naive Mean | 39879.84 / 168.27 / 36.44 |
+| Naive Seasonal | 10.55 / 1.32 / 1.31 | Naive Seasonal | 5057.47 / 52.81 / 11.89 |
+| Naive Drift | 10.60 / 1.15 / 1.30 | Naive Drift | 5466.23 / 60.58 / 14.70 |
+| Naive Moving Average | 12.13 / 1.27 / 1.34 | Naive Moving Average | 5057.47 / 52.81 / 11.89 |
+| N-Beats | 72.11 / 7.10 / 7.40 | N-Beats | 4532.84 / 39.21 / 23.49 |
+| DeepAR | 286.82 / 15.67 / 16.36 | DeepAR | 6325.75 / 35.97 / 16.59 |
+| Prophet | 88.93 / 10.97 / 6.54 | Prophet | 3768.15 / 29.36 / 24.01 |
+| LLMTime with GPT-3.5-Turbo | 76.83 / 7.15 / 7.42 | LLMTime with GPT-3.5-Turbo | 224.54 / 3.07 / 0.83 |
+| LLMTime with GPT-4-Turbo | 988.14 / 26.57 / 29.22 | LLMTime with GPT-4-Turbo | 111.65 / 2.40 / 0.64 |
+| LLMTime with Gemini-1.0-pro | 57.96 / 6.01 / 6.66 | LLMTime with Gemini-1.0-pro | 176.32 / 3.72 / 0.75 |
+| LLMtime with Llama-2 | 75.58 / 7.11 / 7.94 | LLMtime with Llama-2 | 215.39 / 4.07 / 1.31 |
 4.1 Implementation Details
@@
 We tokenize the input sequence and let the LLMs output the period directly.
-We use GPT-3.5-turbo, GPT-4-turbo and Gemini-1.0-Pro to predict the periods.
+We use GPT-3.5-turbo, GPT-4-turbo, and Gemini-1.0-Pro to predict the periods.
 We have chosen five datasets with their seasonal strengths exceeding 85%.
@@
 The mode is selected as the evaluation metric because, when considering the usage characteristics of LLMs, the output of this number best represents the model’s normal performance.
-We defer our details of the prompt in subsection A.3, and the results are shown in Table 2.
+The results are shown in Table 3.
 4.2 Key Findings
@@
 Both methods achieve substantially improved model performance.
+5.1 Dataset description and the External Knowledge incorporated in the Prompts
+We briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts.
+Following [5], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets.
+Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods.
+Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
 5.2 External Knowledge Enhancing Time Series Forecasting
@@
 The core idea of this part is to use the knowledge obtained from the pre-training stage to help predict.
-We provide the large language model with some basic information about the current dataset such as the background of the data collection, and this process does not involve data leakage.
-We incorporate our tests on the data leakage in Appendix A.4.
-It is noted that we do not provide the LLMs with any statistical information such as the periods or trends.
+We provide the large language model with some basic information about the current dataset, such as the background of the data collection, and this process does not involve data leakage.
+We incorporate our tests on the data leakage in Section 5.5.
+We do not provide the LLMs with any statistical information, such as the periods or trends.
 This approach ensures that the LLMs forecast the time series entirely based on the data and their prior knowledge.
@@
 We input the dataset’s external knowledge through prompts before the sequence’s input.
-The external knowledge of each dataset is presented in subsection A.1.
+The external knowledge of each dataset is presented in subsection 5.1.
 The results are shown in Table 8, where LLMTime Prediction refers to the approach described by [5] without any modifications.
@@
 Using External Knowledge Enhancing, Gemini outperforms other models on MonthlyMilk, Sunspots, Wooly, and HeartRate Datasets, but performs poorly on other datasets.
-5.2 Natural Language Paraphrasing
+5.3 Natural Language Paraphrasing Time Series Forecasting
 In this subsection, we conduct experiments on the natural language paraphrasing of the input time sequences.
@@
 We use natural language to describe the trend between consecutive values.
-For instance, given a time series where , we describe the trend from to as follows: "The value rises from to , and falls from to …".
+For instance, given a time series where , we describe the trend from to as follows: ”The value rises from to , and falls from to …”.
 The string we get here is our natural language paraphrasing sequence.
@@
 5.3.1 Implementation Details
+We use GPT-3.5-Turbo, GPT-4-turbo, Llama-2 and Gemini-Pro-1.0 to forecast the time series, where part of the results are presented in Table 4 due to the page limit.
 5.3.2 Key Findings
@@
 All these results demonstrate the superior performance of our methods.
+5.4 Computational Cost
+For reference, we list the average token length cost associated with external knowledge enhancement and natural language paraphrasing.
+Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy.
+It is noted that Natural Language Paraphrasing is judged one by one through hard coding.
+Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time.
+The results are shown in Table 7.
+Several key observations can be made: the original TimeLLM method maintains a uniform token length of 200 across all datasets, providing a stable baseline.
+EKE results in a slight increase in token length, ranging from 7% to 12%, suggesting a good balance between incorporating additional context and maintaining computational efficiency.
+In contrast, NLP leads to a more substantial increase in token numbers.
+5.5 Tests on Data Leakage
+We indirectly explore the data leakage problem by asking LLMs if they can identify the dataset name, the first 20 steps of the predicted dataset, and identify the dataset based on the first 20 steps of the time series data points.
+The results show that although GPT and Gemini can identify and determine data sets with limited information, they generally do not have detailed sequence data knowledge for a wider range of data sets Table 6.
 | Models | Datasets | Natural Language Paraphrasing | LLMTime Prediction | ||||
@@
 | ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
+| Models | Dataset | External Knowledge Enhancing | LLMTime Prediction | ||||
+| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
+| AirPassengers | 3713.99 | 50.37 | 10.88 | 6244.07 | 61.39 | 14.43 | |
+| AusBeer | 669.01 | 21.82 | 5.12 | 841.68 | 23.59 | 5.62 | |
+| GasRateCO2 | 16.47 | 3.36 | 5.97 | 10.88 | 2.66 | 4.73 | |
+| MonthlyMilk | 4781.26 | 55.45 | 6.25 | 7507.13 | 66.28 | 112.77 | |
+| GPT-3.5-turbo-1106 | Sunspots | 7072.42 | 62.61 | 194.29 | 6556.55 | 58.95 | 217.94 |
+| HeartRate | 59.83 | 6.44 | 6.75 | 76.83 | 7.15 | 7.42 | |
+| Istanbul-Traffic | 888.31 | 28.16 | 60.11 | 1321.44 | 48.7 | 7.47 | |
+| ETTh1 | 2.65 | 1.01 | 132.13 | 5.64 | 2.71 | 1.625 | |
+| ETTm2 | 2.00 | 0.89 | 201.84 | 3.46 | 2.17 | 1.178 | |
+| AirPassengers | 1262.24 | 30.54 | 6.80 | 1286.25 | 28.04 | 6.07 | |
+| AusBeer | 345.59 | 15.70 | 3.69 | 513.49 | 18.57 | 4.28 | |
+| GasRateCO2 | 6.99 | 2.29 | 4.21 | 7.27 | 2.32 | 4.18 | |
+| MonthlyMilk | 2209.33 | 44.02 | 5.12 | 4442.18 | 50.75 | 172.82 | |
+| GPT-4-turbo-preview | Sunspots | 4571.92 | 50.24 | 334.30 | 3374.70 | 41.87 | 321.11 |
+| HeartRate | 78.99 | 6.96 | 7.90 | 988.14 | 26.57 | 29.22 | |
+| Istanbul-Traffic | 954.88 | 26.92 | 47.29 | 1291.17 | 32.16 | 6.46 | |
+| ETTh1 | 2.70 | 1.06 | 129.99 | 4.73 | 1.53 | 3.282 | |
+| ETTm2 | 1.18 | 0.79 | 291.67 | 2.30 | 1.034 | 1.607 | |
+| AirPassengers | 3713.99 | 50.37 | 10.88 | 1286.25 | 28.04 | 6.07 | |
+| AusBeer | 893.56 | 21.49 | 4.87 | 644.82 | 17.88 | 4.08 | |
+| GasRateCO2 | 11.38 | 3.04 | 5.49 | 12.78 | 2.97 | 5.47 | |
+| MonthlyMilk | 4722.32 | 60.36 | 7.05 | 3410.20 | 41.40 | 240.25 | |
+| Llama-2 | Sunspots | 4000.19 | 46.45 | 138.69 | 4467.67 | 48.95 | 91.79 |
+| HeartRate | 112.17 | 7.86 | 8.93 | 75.58 | 7.11 | 7.94 | |
+| Istanbul-Traffic | 979.15 | 26.70 | 45.57 | 1531.37 | 34.74 | 7.42 | |
+| ETTh1 | 4.15 | 1.65 | 408.11 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 3.08 | 1.47 | 810.56 | 3.31 | 2.07 | 2.153 | |
+| AirPassengers | 5237.85 | 51.92 | 11.08 | 6392.21 | 63.57 | 14.03 | |
+| AusBeer | 325.45 | 10.84 | 1.86 | 397.78 | 14.36 | 3.27 | |
+| GasRateCO2 | 15.54 | 3.23 | 4.43 | 18.99 | 3.57 | 6.46 | |
+| MonthlyMilk | 491.26 | 15.18 | 1.13 | 628.98 | 17.01 | 1.99 | |
+| Gemini-1.0-pro | Sunspots | 491.64 | 11.15 | 1.27 | 626.03 | 14.94 | 1.73 |
+| HeartRate | 47.45 | 4.83 | 4.67 | 57.96 | 6.01 | 6.66 | |
+| Istanbul-Traffic | 1253.74 | 28.25 | 5.42 | 1531.37 | 34.74 | 7.42 | |
+| ETTh1 | 2.92 | 1.45 | 2.88 | 4.84 | 1.79 | 3.178 | |
+| ETTm2 | 2.00 | 1.74 | 1.22 | 3.31 | 2.07 | 2.153 | |
 6 Related Work
@@
 6.1 Traditional Time Series Forecasting
-Two commonly used methods for traditional time series analysis are the ARIMA method [2] and the exponential smoothing method [7].
+Two commonly used traditional time series analysis methods are the ARIMA method [16] and the exponential smoothing method [17].
 The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions.
@@
 The ARIMA model requires testing data stationarity and selecting the right order.
-However, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower than the ARIMA model.
+However, the exponential smoothing method is not affected by outliers; it is only suitable for stationary time series, and its accuracy in predicting future values is lower than that of the ARIMA model.
 6.2 LLMs for Time Series Forecasting
-The first family of methods involve either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [17, 8, 6, 3].
+The first family of methods involves either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [6, 18, 19, 20].
 For instance, [6] aimed to build the foundational models for time series and investigate its scaling behavior.
@@
 Although these studies contribute significantly to understanding foundational models, they require considerable computing resources and expertise in fine-tuning procedures.
-Moreover, the details of the model may not be disclosed for commercial purposes [8], which impedes future research.
+Moreover, the details of the model may not be disclosed for commercial purposes [18, 22], which impedes future research.
 Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
 In contrast, the second family of methods does not involve model parameter finetuning.
-These methods either create appropriate prompts or reprogramme inputs, to effectively handle time series data [10, 19, 14, 21].
+These methods either create appropriate prompts or reprogramme inputs to effectively handle time series data [5, 7, 23, 24].
 [7] tokenizes the time series and manages to embed those tokens, and [23] reprogrammed the time series data with text prototypes before feeding them to the LLMs.
@@
 7 Conclusions and Future Work
-In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero shot setting, revealing a proclivity for data with distinct trends and seasonal patterns.
+In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero-shot setting, revealing a proclivity for data with distinct trends and seasonal patterns.
 Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity.
@@
 Lastly, experimental results indicate that our proposed strategies of incorporating external knowledge and transforming numerical sequences into natural language formats have yielded substantial improvements in accuracy.
-Limitation
-This study may be limited in the following ways.
-First, limitations in the scope of the dataset and large language models may not capture the full variability of the results with a wider array.
-In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods.
-Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the model’s performance in different data domains.
-These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, underscoring the need to expand future research.
-Appendix A Appendix
-A.1 Dataset description and the External Knowledge incorporated in the Prompts
-In this subsection, we briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts.
-Following [10], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets.
-Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods.
-Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
-A.1.1 Darts [11]
-(1)AirPassengersDataset: This is a series of monthly passenger numbers for international flights, where each value is in thousands of passengers for that month.
-(2)AusBeerDataset: This is a quarterly series of beer production, with each value representing the kiloliters of beer produced in that quarter.
-(3)GasRateCO2Dataset: This time series dataset describes monthly carbon dioxide emissions.
-(4)MonthlyMilkDataset: This time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
-(5)SunspotsDataset: This dataset records the number of sunspots each month, where each data is the number of sunspots in that month.
-(6)WineDataset: This is a dataset of monthly wine production in Australia, where each figure is the number of wine bottles produced in that month.
-(7)WoolyDataset: This is an Australian yarn production for each quarter, where each value is how many tons of yarn were produced in that quarter.
-(8)HeartRateDataset: The series contains 1800 uniformly spaced instantaneous heart rate measurements from a single subject.
-(9)ETTh1: This is a time series dataset containing high-frequency energy data of a certain region in China, which is mainly used for energy load forecasting and related time series analysis research.
-(10)ETTm2: Similar to ETTh1, ETTm2 is also an energy time series dataset, but the data frequency or region covered may be different, which is also used for the analysis and prediction of energy consumption.
-A.1.2 Memorization Datasets [10]
-(11)TSMCStockDataset: This is historical trading data about Taiwan Semiconductor Manufacturing Corporation (TSMC) stock, containing information such as share price, volume, and date, and is commonly used in financial analysis and stock market forecasting research.
-(12)TurkeyPowerDataset: This is a time series dataset on national electricity consumption in Turkey, which records the electricity usage in Turkey over a period of time and is often used to analyze and forecast electricity demand.
-(13)IstanbulTrafficDataset: This dataset offers hourly Traffic Index data for Istanbul from October 2022 to May 2023.
-A.1.3 Monash Datasets [9]
-(14)US Births Dataset: This dataset contains the number of births in the US from 1969 to 1988.
-(15)Saugeen River Flow Dataset: This dataset contains the daily mean flow of the Saugeen River at Walkerton in cubic meters per second from 1915 to 1979.
-A.2 Implementation Details
-In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data.
-We first describe both the real and synthesized datasets we use and then detail the methods we use to investigate the preferences of LLMs.
-A.2.1 Real Datasets
-We begin by comparing the performance of Large Language Models (LLMs) with traditional time series forecasting methods.
-The results are depicted in Figures Figure 4 and Figure 5, while Tables Table 9 and Table 10 present the computed metrics in tabular format.
-To understand the preferences of LLMs, we conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTrafficDataset and TurkeyPowerDataset.
-We apply the Seasonal-Trend decomposition using the LOESS (STL) technique to decompose the original time series into trend, seasonal, and residual components.
-In those datasets, we obtain the periods through the nature of the data.
-For instance, the number of passengers is collected monthly in AirPassengersDataset, and it’s natural to obtain that the period is 12.
-For the datasets without explicit periods, such as the IstanbulTrafficDataset, the period is determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period.
-The strengths and the model performance can be seen in Table 6.
-Subsequently, we compute trend strength and seasonal strength to measure all of those components.
-We use MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 1).
-Notably, GPT-4-turbo achieved a higher absolute PCC compared to GPT-3.5-turbo.
-It may be attributed to human feedback during GPT-4-turbo training, as individuals may be more aware of seasonal and trend data.
-This may provide some insights for further research into the characteristics of the LLMs with time series forecasting.
-We also conduct a counterfactual analysis using a systematic permutation of the input time series.
-We begin by defining a sliding window that determines the length of the periods and add Gaussian noise to the data within this window.
-To reduce costs, we move the window by the period length.
-This method allows us to assess the importance of segments that LLM predominantly focuses on.
-Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
-A.2.2 Synthesized Datasets
-To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components.
-We set , , and , uniformly sampled for 10 instances each, and and to represent the chosen frequencies.
-Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution .
-Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure 8.
-This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
-A.3 Detailed Prompts for the Period Prediction
-GPT-3.5-turbo and GPT-4-turbo.
-You are a helpful assistant that specializes in time series analysis.
-The user will provide a sequence.
-The sequence is separated by commas.
-You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence.
-Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence.
-The underlying cycle of the sequence is strictly below 8.
-Please infer the cycle without producing any additional text.
-Again, the sequence is separated by commas.
-Sequence:
-Gemini.
-You are a helpful assistant that specializes in time series analysis.
-The user will provide a sequence.
-The sequence is represented by decimal strings separated by commas.
-You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence.
-Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence.
-The underlying cycle of the sequence is strictly below 8.
-Please infer the cycle without producing any additional text.
-Sequence:
-A.4 Tests on Data Leakage
-Since the detailed datasets used for training GPT-4 and Gemini are not directly accessible, we conducted several experiments to indirectly investigate the data leakage issue.
-First, we performed an "Acknowledge Test" where we asked the LLMs if they recognize the dataset based on its name.
-Next, we conducted a "Series Test" by asking the LLMs to predict the first 20 steps of the datasets.
-Finally, we carried out a "Dataset Detection" test, where we fed the first 20 steps into the LLMs and asked them to identify the dataset.
-All these experiments help us determine whether the LLMs only recognize the name or have detailed knowledge of the datasets.
-As shown in 4, we observe that both GPT and Gemini are generally aware of the content of most datasets.
-However, detailed sequence data is primarily known only for the AirPassengers dataset when using GPT, and this level of detail is not known for other datasets.
-During the experiments on the Dataset Detection, it is found that both GPT and Gemini could identify which dataset is being used based on the first 20 data points of the time series.
-All these results indicate that while they can recognize and identify datasets from limited information, they typically do not possess detailed knowledge of the sequence data for a broader range of datasets.
 | Datasets | Acknowledge Test (GPT) | Acknowledge Test (Gemini) | Series Test (GPT) | Series Test (Gemini) | Dataset Detection (GPT) | Dataset Detection (Gemini) |
@@
 | HeartRate | Yes | Yes | No | No | No | No |
-A.5 Computational Cost
-We list the average token length cost associated with external knowledge enhancing and natural language paraphrasing for reference.
-Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy.
-It is noted that Natural Language Paraphrasing is judged one by one through hard coding.
-Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time.
-The results are shown in 5.
 | Datasets | Avg Token Length (ori) | Avg Token Length (EKE) | Avg Token Length (NLP) |
@@
 | Dataset Name | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
-| AirPassengersDataset | 6.80 | 9.98 | 1.00 | 0.98 |
-| AusBeerDataset | 3.69 | 5.12 | 0.99 | 0.96 |
-| MonthlyMilkDataset | 5.12 | 6.25 | 1.00 | 0.99 |
-| SunspotsDataset | 334.30 | 194.29 | 0.81 | 0.28 |
-| WineDataset | 10.90 | 14.98 | 0.67 | 0.92 |
-| WoolyDataset | 20.41 | 19.26 | 0.96 | 0.82 |
-| IstanbulTrafficGPT | 47.29 | 60.11 | 0.31 | 0.72 |
-| GasRateCO2Dataset | 4.21 | 5.97 | 0.65 | 0.50 |
-| HeartRateDataset | 7.90 | 6.75 | 0.42 | 0.49 |
-| TurkeyPower | 3.36 | 3.52 | 0.90 | 0.88 |
-A.6 Figures and Tables
-The predicted results of AirPassengers, AusBeerDataset, GasRateCO2, HeartRate, Istanbul-Traffic datasets.
-The predicted results of MonthlyMilk, Sunspots, TSMCStock, TurkeyPower, WineDataset, Wooly datasets.
-| Models | Datasets | Natural Language Paraphrasing | LLMTime Prediction | ||||
-| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
-| GPT-3.5-Turbo | AirPassengers | 267.66 | 3.66 | 0.99 | 6244.07 | 61.39 | 14.43 |
-| AusBeer | 598.45 | 5.81 | 1.36 | 841.68 | 23.59 | 5.62 | |
-| GasRateCO2 | 3.16 | 0.46 | 0.85 | 10.88 | 2.66 | 4.73 | |
-| MonthlyMilk | 968.69 | 8.61 | 1.02 | 7507.13 | 66.28 | 112.77 | |
-| Sunspots | 251.61 | 4.27 | 20.42 | 6556.55 | 58.95 | 217.94 | |
-| (GPT-3.5-turbo-1106) | Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 |
-| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
-| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
-| Istanbul-Traffic | 224.17 | 3.74 | 8.81 | 335.05 | 6.75 | 11.68 | |
-| Turkey Power | 24382136.98 | 1843.64 | 4.68 | 3882704.14 | 1315.6 | 3.58 | |
-| ETTh1 | 1.21 | 0.48 | 54.17 | 5.64 | 2.71 | 1.625 | |
-| ETTm2 | 0.81 | 0.36 | 27.33 | 3.46 | 2.17 | 1.178 | |
-| US Births | 926136.72 | 633.37 | 7.67 | 1323052.46 | 904.81 | 9.61 | |
-| Saugeen River Flow | 2870.19 | 18.31 | 19.43 | 4100.27 | 26.16 | 27.76 | |
-| GPT-4-Turbo | AirPassengers | 133.10 | 2.87 | 0.80 | 1286.25 | 28.04 | 6.07 |
-| AusBeer | 661.80 | 7.24 | 1.63 | 513.49 | 18.57 | 4.28 | |
-| GasRateCO2 | 2.28 | 0.41 | 0.75 | 7.27 | 2.32 | 4.18 | |
-| MonthlyMilk | 413.63 | 4.94 | 0.57 | 4442.18 | 50.75 | 172.82 | |
-| Sunspots | 194.52 | 5.30 | 16.10 | 3374.70 | 41.87 | 321.11 | |
-| (GPT-4-turbo-preview) | Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 |
-| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
-| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
-| Istanbul-Traffic | 176.91 | 3.88 | 9.67 | 195.33 | 5.53 | 10.03 | |
-| Turkey Power | 60601807.53 | 3118.43 | 8.03 | 113873.28 | 814.46 | 2.17 | |
-| ETTh1 | 1.20 | 0.49 | 47.62 | 4.73 | 1.53 | 3.282 | |
-| ETTm2 | 0.45 | 0.27 | 23.62 | 2.30 | 1.034 | 1.607 | |
-| US Births | 676264.81 | 501.13 | 4.81 | 966092.59 | 678.55 | 7.12 | |
-| Saugeen River Flow | 3100.57 | 19.18 | 21.86 | 4190.05 | 27.41 | 32.15 | |
-| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1317.9 | 55.49 | 11.18 |
-| AusBeer | 591.75 | 23.25 | 5.41 | 644.82 | 17.88 | 4.08 | |
-| GasRateCO2 | 10.16 | 2.89 | 5.16 | 12.78 | 2.97 | 5.47 | |
-| MonthlyMilk | 851.17 | 84.83 | 9.46 | 3410.20 | 41.40 | 240.25 | |
-| Sunspots | 1483.29 | 33.27 | 17.79 | 4467.67 | 48.95 | 91.79 | |
-| (llama-2-13B) | Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 |
-| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
-| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
-| Istanbul-Traffic | 306.80 | 5.39 | 7.24 | 438.28 | 7.28 | 9.81 | |
-| Turkey Power | 3278744.18 | 2191.34 | 28.76 | 2919773.15 | 1388.10 | 3.70 | |
-| ETTh1 | 1.47 | 0.87 | 58.34 | 4.84 | 1.79 | 3.178 | |
-| ETTm2 | 0.84 | 0.41 | 29.86 | 3.31 | 2.07 | 2.153 | |
-| US Births | - | - | - | - | - | - | |
-| Saugeen River Flow | - | - | - | - | - | - | |
-| Gemini-Pro-1.0 | AirPassengers | 4474.54 | 31.54 | 7.02 | 6392.21 | 63.57 | 14.03 |
-| AusBeer | 278.45 | 10.05 | 2.29 | 397.78 | 14.36 | 3.27 | |
-| GasRateCO2 | 13.29 | 2.50 | 4.38 | 18.99 | 3.57 | 6.46 | |
-| MonthlyMilk | 440.29 | 11.91 | 1.39 | 628.98 | 17.01 | 1.99 | |
-| Sunspots | 438.29 | 10.47 | 1.21 | 626.03 | 14.94 | 1.73 | |
-| (gemini-1.0-pro) | Wine | 181008.34 | 2557.62 | 10.89 | 258584.78 | 3645.23 | 14.60 |
-| Wooly | 45.44 | 4.47 | 4.93 | 64.92 | 6.39 | 7.04 | |
-| HeartRate | 40.57 | 4.20 | 4.67 | 57.96 | 6.01 | 6.66 | |
-| Istanbul-Traffic | 267.43 | 5.69 | 8.37 | 321.56 | 7.32 | 9.71 | |
-| Turkey Power | 45674.41 | 2973.54 | 11.21 | 103203.37 | 2195.68 | 6.07 | |
-| ETTh1 | 1.17 | 0.74 | 54.86 | 4.84 | 1.79 | 3.178 | |
-| ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
-| US Births | 467743.19 | 440.68 | 4.27 | 687862.05 | 595.51 | 6.28 | |
-| Saugeen River Flow | 2857.23 | 19.34 | 18.80 | 4081.75 | 25.91 | 26.87 | |
-| Models | Dataset | External Knowledge Enhancing | LLMTime Prediction | ||||
-| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
-| AirPassengers | 3713.99 | 50.37 | 10.88 | 6244.07 | 61.39 | 14.43 | |
-| AusBeer | 669.01 | 21.82 | 5.12 | 841.68 | 23.59 | 5.62 | |
-| GasRateCO2 | 16.47 | 3.36 | 5.97 | 10.88 | 2.66 | 4.73 | |
-| MonthlyMilk | 4781.26 | 55.45 | 6.25 | 7507.13 | 66.28 | 112.77 | |
-| Sunspots | 7072.42 | 62.61 | 194.29 | 6556.55 | 58.95 | 217.94 | |
-| GPT-3.5-turbo-1106 | Wine | 24925885.81 | 3548.19 | 14.98 | 30488.60 | 388.28 | 15.83 |
-| (GPT-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 526903.08 | 574.58 | 12.00 |
-| HeartRate | 59.83 | 6.44 | 6.75 | 76.83 | 7.15 | 7.42 | |
-| Istanbul-Traffic | 888.31 | 28.16 | 60.11 | 1321.44 | 48.7 | 7.47 | |
-| TSMC-Stock | 73.83 | 7.31 | 1.54 | 298.58 | 15.44 | 3.23 | |
-| Turkey Power | 2613198.17 | 1301.83 | 3.52 | 3882704.14 | 1315.6 | 3.58 | |
-| ETTh1 | 2.65 | 1.01 | 132.13 | 5.64 | 2.71 | 1.625 | |
-| ETTm2 | 2.00 | 0.89 | 201.84 | 3.46 | 2.17 | 1.178 | |
-| AirPassengers | 1262.24 | 30.54 | 6.80 | 1286.25 | 28.04 | 6.07 | |
-| AusBeer | 345.59 | 15.70 | 3.69 | 513.49 | 18.57 | 4.28 | |
-| GasRateCO2 | 6.99 | 2.29 | 4.21 | 7.27 | 2.32 | 4.18 | |
-| MonthlyMilk | 2209.33 | 44.02 | 5.12 | 4442.18 | 50.75 | 172.82 | |
-| Sunspots | 4571.92 | 50.24 | 334.30 | 3374.70 | 41.87 | 321.11 | |
-| GPT-4-turbo-preview | Wine | 14426570.88 | 2734.41 | 10.90 | 22488.17 | 253.08 | 9.98 |
-| (GPT-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 942987.19 | 871.64 | 18.55 |
-| HeartRate | 78.99 | 6.96 | 7.90 | 988.14 | 26.57 | 29.22 | |
-| Istanbul-Traffic | 954.88 | 26.92 | 47.29 | 1291.17 | 32.16 | 6.46 | |
-| TSMC-Stock | 104.53 | 8.46 | 1.79 | 74.71 | 6.60 | 1.39 | |
-| Turkey Power | 3090055.89 | 1223.78 | 3.36 | 113873.28 | 814.46 | 2.17 | |
-| ETTh1 | 2.70 | 1.06 | 129.99 | 4.73 | 1.53 | 3.282 | |
-| ETTm2 | 1.18 | 0.79 | 291.67 | 2.30 | 1.034 | 1.607 | |
-| AirPassengers | 3713.99 | 50.37 | 10.88 | 1286.25 | 28.04 | 6.07 | |
-| AusBeer | 893.56 | 21.49 | 4.87 | 644.82 | 17.88 | 4.08 | |
-| GasRateCO2 | 11.38 | 3.04 | 5.49 | 12.78 | 2.97 | 5.47 | |
-| MonthlyMilk | 4722.32 | 60.36 | 7.05 | 3410.20 | 41.40 | 240.25 | |
-| Sunspots | 4000.19 | 46.45 | 138.69 | 4467.67 | 48.95 | 91.79 | |
-| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 951194.94 | 240.08 | 9.45 |
-| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 675062.52 | 736.04 | 15.83 |
-| HeartRate | 112.17 | 7.86 | 8.93 | 75.58 | 7.11 | 7.94 | |
-| Istanbul-Traffic | 979.15 | 26.70 | 45.57 | 1531.37 | 34.74 | 7.42 | |
-| TSMC-Stock | 52105.36 | 196.02 | 42.07 | 2203.97 | 27.64 | 27.39 | |
-| Turkey Power | 3416162.71 | 1547.49 | 4.09 | 2919773.15 | 1388.10 | 3.70 | |
-| ETTh1 | 4.15 | 1.65 | 408.11 | 4.84 | 1.79 | 3.178 | |
-| ETTm2 | 3.08 | 1.47 | 810.56 | 3.31 | 2.07 | 2.153 | |
-| AirPassengers | 5237.85 | 51.92 | 11.08 | 6392.21 | 63.57 | 14.03 | |
-| AusBeer | 325.45 | 10.84 | 1.86 | 397.78 | 14.36 | 3.27 | |
-| GasRateCO2 | 15.54 | 3.23 | 4.43 | 18.99 | 3.57 | 6.46 | |
-| MonthlyMilk | 491.26 | 15.18 | 1.13 | 628.98 | 17.01 | 1.99 | |
-| Sunspots | 491.64 | 11.15 | 1.27 | 626.03 | 14.94 | 1.73 | |
-| Gemini-1.0-pro | Wine | 210818.24 | 3230.41 | 8.35 | 258584.78 | 3645.23 | 14.60 |
-| (gemini-1.0-pro) | Wooly | 51.04 | 5.70 | 7.93 | 64.92 | 6.39 | 7.04 |
-| HeartRate | 47.45 | 4.83 | 4.67 | 57.96 | 6.01 | 6.66 | |
-| Istanbul-Traffic | 1253.74 | 28.25 | 5.42 | 1531.37 | 34.74 | 7.42 | |
-| TSMC-Stock | 153.73 | 5.02 | 1.05 | 188.18 | 6.67 | 1.65 | |
-| Turkey Power | 83522.95 | 1812.31 | 5.51 | 103203.37 | 2195.68 | 6.07 | |
-| ETTh1 | 2.92 | 1.45 | 2.88 | 4.84 | 1.79 | 3.178 | |
-| ETTm2 | 2.00 | 1.74 | 1.22 | 3.31 | 2.07 | 2.153 | |
-| Dataset | Method | MSE | MAE | MAPE |
-| AirPassengers | Exponential Smoothing | 2007.67 | 37.91 | 8.10 |
-| SARIMA | 2320.47 | 39.80 | 8.46 | |
-| Cyclical Regression | 2028.37 | 36.70 | 8.52 | |
-| AutoARIMA | 8702.09 | 68.52 | 13.98 | |
-| FFT | 3274.46 | 46.38 | 10.59 | |
-| StatsForecastAutoARIMA | 2952.52 | 45.41 | 9.71 | |
-| Naive Mean | 47703.65 | 204.25 | 44.61 | |
-| Naive Seasonal | 6032.80 | 62.87 | 14.18 | |
-| Naive Drift | 6505.79 | 72.21 | 17.50 | |
-| Naive Moving Average | 6032.80 | 62.87 | 14.18 | |
-| N-Beats | 3994.55 | 54.95 | 12.81 | |
-| DeepAR | 184222.64 | 421.99 | 98.42 | |
-| Prophet | 7345.31 | 43.87 | 8.62 | |
-| LLMTime with GPT-3.5-Turbo | 6244.07 | 61.39 | 14.43 | |
-| LLMTime with GPT-4-Turbo | 1317.9 | 55.49 | 11.18 | |
-| LLMTime with Gemini-1.0-pro | 6392.21 | 63.57 | 14.03 | |
-| LLMtime with Llama-2 | 1286.25 | 28.04 | 6.07 | |
-| AusBeer | Exponential Smoothing | 703.26 | 22.80 | 5.44 |
-| SARIMA | 475.53 | 19.07 | 4.49 | |
-| Cyclical Regression | 989.31 | 26.29 | 6.13 | |
-| AutoARIMA | 550.05 | 18.84 | 4.41 | |
-| FFT | 7682.56 | 73.74 | 17.44 | |
-| StatsForecastAutoARIMA | 559.46 | 20.56 | 4.86 | |
-| Naive Mean | 1885.72 | 30.66 | 6.68 | |
-| Naive Seasonal | 10828.02 | 96.35 | 23.39 | |
-| Naive Drift | 18507.61 | 128.23 | 30.91 | |
-| Naive Moving Average | 10828.02 | 96.35 | 23.39 | |
-| N-Beats | 250.61 | 14.42 | 3.53 | |
-| DeepAR | 16197.17 | 40.23 | 9.89 | |
-| Prophet | 6323.89 | 28.76 | 6.92 | |
-| LLMTime with GPT-3.5-Turbo | 841.68 | 23.59 | 5.62 | |
-| LLMTime with GPT-4-Turbo | 513.49 | 18.57 | 4.28 | |
-| LLMTime with Gemini-1.0-pro | 397.78 | 14.36 | 3.27 | |
-| LLMtime with Llama-2 | 644.82 | 17.88 | 4.08 | |
-| MonthlyMilk | Exponential Smoothing | 564.94 | 20.23 | 2.41 |
-| SARIMA | 1289.76 | 32.78 | 3.87 | |
-| Cyclical Regression | 3631.53 | 56.15 | 6.60 | |
-| AutoARIMA | 2682.67 | 42.82 | 5.20 | |
-| FFT | 3453.96 | 45.62 | 5.48 | |
-| StatsForecastAutoARIMA | 186.14 | 10.64 | 1.28 | |
-| Naive Mean | 19893.07 | 127.33 | 14.46 | |
-| Naive Seasonal | 4870.40 | 56.00 | 6.31 | |
-| Naive Drift | 3998.11 | 56.06 | 6.52 | |
-| Naive Moving Average | 4870.40 | 56.00 | 6.31 | |
-| N-Beats | 3140.89 | 51.57 | 6.07 | |
-| DeepAR | 728289.50 | 851.30 | 99.22 | |
-| Prophet | 663.41 | 25.76 | 2.92 | |
-| LLMTime with GPT-3.5-Turbo | 7507.13 | 66.28 | 112.77 | |
-| LLMTime with GPT-4-Turbo | 4442.18 | 50.75 | 172.82 | |
-| LLMTime with Gemini-1.0-pro | 628.98 | 17.01 | 1.99 | |
-| LLMtime with Llama-2 | 3410.20 | 41.40 | 240.25 | |
-| Sunspots | Moving Average | 326750.49 | 499.78 | 3129.63 |
-| Exponential Smoothing | 326750.49 | 499.78 | 3129.63 | |
-| SARIMA | 2902.72 | 45.75 | 466.99 | |
-| Cyclical Regression | 3917.76 | 47.84 | 274.31 | |
-| AutoARIMA | 4695.67 | 58.47 | 709.23 | |
-| FFT | 3784.56 | 49.81 | 150.32 | |
-| StatsForecastAutoARIMA | 8406.55 | 72.99 | 95.18 | |
-| Naive Mean | 4120.40 | 49.84 | 267.22 | |
-| Naive Seasonal | 4440.63 | 56.78 | 688.58 | |
-| Naive Drift | 5032.77 | 60.40 | 724.88 | |
-| Naive Moving Average | 4440.63 | 56.78 | 688.58 | |
-| N-Beats | 4877.59 | 56.58 | 105.55 | |
-| DeepAR | 3421.02 | 48.93 | 132.76 | |
-| Prophet | 6303.57 | 76.83 | 67.97 | |
-| LLMTime with GPT-3.5-Turbo | 6556.55 | 58.95 | 217.94 | |
-| LLMTime with GPT-4-Turbo | 3374.70 | 41.87 | 321.11 | |
-| LLMTime with Gemini-1.0-pro | 626.03 | 14.94 | 1.73 | |
-| LLMtime with Llama-2 | 4467.67 | 48.95 | 91.79 |
-| Dataset | Method | MSE | MAE | MAPE |
-| WineDataset | Exponential Smoothing | 23709576.52 | 3370.78 | 14.23 |
-| SARIMA | 1150166.94 | 966.57 | 20.76 | |
-| Cyclical Regression | 7873785.27 | 2148.24 | 8.52 | |
-| AutoARIMA | 698661.90 | 646.03 | 14.07 | |
-| FFT | 1031170.45 | 867.83 | 18.60 | |
-| StatsForecastAutoARIMA | 20040877.37 | 2853.17 | 12.05 | |
-| Naive Mean | 11557786.19 | 2200.04 | 8.80 | |
-| Naive Seasonal | 879447.22 | 724.23 | 15.52 | |
-| Naive Drift | 9609576.04 | 1833.38 | 7.36 | |
-| Naive Moving Average | 9070696.99 | 1719.17 | 6.90 | |
-| N-Beats | 5418377.00 | 1887.30 | 7.68 | |
-| DeepAR | 715027008.00 | 26236.14 | 89.91 | |
-| Prophet | 4846922.27 | 2201.57 | 8.27 | |
-| LLMTime with GPT-3.5-Turbo | 30488.60 | 388.28 | 15.83 | |
-| LLMTime with GPT-4-Turbo | 22488.17 | 253.08 | 9.98 | |
-| LLMTime with Gemini-1.0-pro | 258584.78 | 3645.23 | 14.60 | |
-| LLMtime with Llama-2 | 951194.94 | 240.08 | 9.45 | |
-| WoolyDataset | Exponential Smoothing | 24925885.81 | 3548.19 | 14.98 |
-| SARIMA | 812352.21 | 759.07 | 16.37 | |
-| Cyclical Regression | 1032574.82 | 962.72 | 22.14 | |
-| AutoARIMA | 838852.91 | 786.25 | 16.84 | |
-| FFT | 1012255.35 | 945.20 | 20.80 | |
-| StatsForecastAutoARIMA | 917617.19 | 858.57 | 18.91 | |
-| Naive Mean | 816762.31 | 764.73 | 16.12 | |
-| Naive Seasonal | 1051110.81 | 982.25 | 22.19 | |
-| Naive Drift | 812352.21 | 759.07 | 16.37 | |
-| Naive Moving Average | 1032574.82 | 962.72 | 22.14 | |
-| N-Beats | 653104.31 | 743.54 | 15.96 | |
-| DeepAR | 243831.14 | 4897.85 | 94.89 | |
-| Prophet | 365241.98 | 891.70 | 34.65 | |
-| LLMTime with GPT-3.5-Turbo | 526903.08 | 574.58 | 12.00 | |
-| LLMTime with GPT-4-Turbo | 942987.19 | 871.64 | 18.55 | |
-| LLMTime with Gemini-1.0-pro | 64.92 | 6.39 | 7.04 | |
-| LLMtime with Llama-2 | 675062.52 | 736.04 | 15.83 | |
-| HeartRateDataset | Exponential Smoothing | 11.16 | 1.38 | 1.49 |
-| SARIMA | 12.98 | 1.34 | 1.61 | |
-| Cyclical Regression | 13.58 | 1.31 | 1.20 | |
-| AutoARIMA | 13.26 | 1.25 | 1.39 | |
-| FFT | 13.95 | 1.16 | 1.34 | |
-| StatsForecastAutoARIMA | 10.53 | 1.27 | 1.39 | |
-| Naive Mean | 12.02 | 1.27 | 1.26 | |
-| Naive Seasonal | 10.55 | 1.32 | 1.31 | |
-| Naive Drift | 10.60 | 1.15 | 1.30 | |
-| Naive Moving Average | 12.13 | 1.27 | 1.34 | |
-| N-Beats | 72.11 | 7.10 | 7.40 | |
-| DeepAR | 286.82 | 15.67 | 16.36 | |
-| Prophet | 88.93 | 10.97 | 6.54 | |
-| LLMTime with GPT-3.5-Turbo | 76.83 | 7.15 | 7.42 | |
-| LLMTime with GPT-4-Turbo | 988.14 | 26.57 | 29.22 | |
-| LLMTime with Gemini-1.0-pro | 57.96 | 6.01 | 6.66 | |
-| LLMtime with Llama-2 | 75.58 | 7.11 | 7.94 | |
-| Weather | Exponential Smoothing | 1684.38 | 31.60 | 6.79 |
-| SARIMA | 1943.81 | 33.33 | 7.09 | |
-| Cyclical Regression | 1700.73 | 30.77 | 7.15 | |
-| AutoARIMA | 7315.10 | 57.44 | 11.70 | |
-| FFT | 2752.02 | 38.90 | 8.87 | |
-| StatsForecastAutoARIMA | 2479.55 | 38.06 | 8.16 | |
-| Naive Mean | 39879.84 | 168.27 | 36.44 | |
-| Naive Seasonal | 5057.47 | 52.81 | 11.89 | |
-| Naive Drift | 5466.23 | 60.58 | 14.70 | |
-| Naive Moving Average | 5057.47 | 52.81 | 11.89 | |
-| N-Beats | 4532.84 | 39.21 | 23.49 | |
-| DeepAR | 6325.75 | 35.97 | 16.59 | |
-| Prophet | 3768.15 | 29.36 | 24.01 | |
-| LLMTime with GPT-3.5-Turbo | 224.54 | 3.07 | 0.83 | |
-| LLMTime with GPT-4-Turbo | 111.65 | 2.40 | 0.64 | |
-| LLMTime with Gemini-1.0-pro | 176.32 | 3.72 | 0.75 | |
-| LLMtime with Llama-2 | 215.39 | 4.07 | 1.31 |
+| AirPassengersDataset | ||||
+| AusBeerDataset | ||||
+| MonthlyMilkDataset | ||||
+| SunspotsDataset | ||||
+| WineDataset | ||||
+| WoolyDataset | ||||
+| IstanbulTrafficGPT | ||||
+| GasRateCO2Dataset | ||||
+| HeartRateDataset | ||||
+| TurkeyPower |